### v2

# references.bib

Abstract
For a graph family $\mathcal{F}$, let $\mathrm{ex}(n,\mathcal{F})$ and $\mathrm{spex}(n,\mathcal{F})$ denote the maximum number of edges and maximum spectral radius of an $n$-vertex $\mathcal{F}$-free graph, respectively, and let $\mathrm{EX}(n,\mathcal{F})$ and $\mathrm{SPEX}(n,\mathcal{F})$ denote the corresponding sets of extremal graphs. Wang, Kang, and Xue showed that if $\mathrm{EX}(n,\mathcal{F})$ consists of Turán graphs $T_{n,r}$ plus $O(1)$ edges, then $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$ for $n$ large enough. Fang, Tait, and Zhai extended this result by showing if $e(T_{n,r})\leq\mathrm{ex}(n,\mathcal{F})<e(T_{n,r})+\lfloor n/2r\rfloor$ then $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$ for $n$ large enough. In this paper we extend the result further and for many values of $r$ we can show that our result is best possible, answering a question of Fang, Tait, and Zhai.
1 Introduction
Let $\mathcal{F}$ be a family of graphs. For $n\in\mathbb{N}$, let $\mathrm{ex}(n,\mathcal{F})$ ($\mathrm{spex}(n,\mathcal{F})$) denote the maximum number of edges (spectral radius of adjacency matrix) of an $n$-vertex graph with no subgraph isomorphic to a graph in $\mathcal{F}$. Let $\mathrm{EX}(n,\mathcal{F})$ ($\mathrm{SPEX}(n,\mathcal{F})$) be the set of graphs attaining the maximum number of edges (spectral radius) among $\mathcal{F}$-free graphs. The problem of determining $\mathrm{ex}(n,\mathcal{F})$ is called the Turán problem and the problem of determining $\mathrm{spex}(n,\mathcal{F})$ is called the spectral Turán problem for $\mathcal{F}$. The latter problem, introduced systematically by Nikiforov [nikiforov2010spectral], has since grown quite popular; for an excellent survey on this topic we refer the reader to [li2021survey]. In particular there have been many papers addressing the question of when we can guarantee that $\mathrm{SPEX}(n,\mathcal{F})=\mathrm{EX}(n,\mathcal{F})$. Some early results in this direction are the spectral Turán theorem [nikiforov2007bounds] which shows that $\mathrm{SPEX}(n,K_{r+1})=\mathrm{EX}(n,K_{r+1})$ and the work on odd cycles [nikiforov2008spectral] which shows that $\mathrm{SPEX}(n,C_{2k+1})=\mathrm{EX}(n,C_{2k+1})$ for large enough $n$. However, this phenomenon does not always occur; for example $\mathrm{SPEX}(n,C_{2k})\cap\mathrm{EX}(n,C_{2k})=\emptyset$ and in fact the extremal families are very different structurally [cioabua2022spectral2, nikiforov2010spectral, zhai2020spectral, zhai2012proof]. Recently Liu and Ning [liu2023spectral] proposed the problem of determining all graphs $F$ for which $\mathrm{SPEX}(n,F)\subseteq\mathrm{EX}(n,F)$ for large enough $n$.
In this paper we ask for a condition on the value of $\mathrm{ex}(n,\mathcal{F})$ which implies some relation between $\mathrm{EX}(n,\mathcal{F})$ and $\mathrm{SPEX}(n,\mathcal{F})$. In [byrne2024general] this question was addressed when $\mathcal{F}$ contains a bipartite graph, so here we focus on the case where all graphs in $\mathcal{F}$ are non-bipartite. The first positive result in this direction was proved by Wang, Kang, and Xue [wang2023conjecture], answering a conjecture of Cioabă, Desai, and Tait [cioabua2022spectral]. Recently, the result of [wang2023conjecture] was extended by Fang, Tait, and Zhai [fang2024turan]. Below, $\chi(\mathcal{F})$ denotes $\min\{\chi(F):F\in\mathcal{F})\}$.
Theorem 1 (Fang, Tait, Zhai [fang2024turan]).
Let $n$ be sufficiently large and let $\mathcal{F}$ be a finite graph family satisfying $\chi(\mathcal{F})=r+1\geq 3$ and $\mathrm{ex}(n,\mathcal{F})<e(T_{n,r})+\left\lfloor\frac{n}{2r}\right\rfloor$. Then $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$.
Motivated by this result, the authors of [fang2024turan] posed the following problem.
Problem 1.
Determine the maximum constant $c=c(r)$ such that for any finite graph family $\mathcal{F}$ with $\chi(\mathcal{F})=r+1\geq 3$, for all positive $\varepsilon<c$ and sufficiently large $n$, $\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,r})+(c-\varepsilon)n$ implies $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$.
Thus, Theorem 1 shows that $c(r)\geq 1/2r$ for every $r$. In this paper we give improved lower and upper bounds on $c(r)$ when $r\geq 3$ and in many cases we can determine $c(r)$ exactly (see 1 below). When $r=2$, we prove that $c(2)\leq 1/3$ and conjecture that this is the correct value. These results also improve the upper bound $c(2)\leq 3/2$ obtained by considering the odd wheel $W_{15}$ [yuan2020extremal, cioabua2022spectral] (we are not aware of existing upper bounds for larger $r$).
The authors of [fang2024turan] also asked the following.
Problem 2.
For $r\geq 2$ and $n$ sufficiently large, if a finite graph family $\mathcal{F}$ satisfies that $e(G)=e(T_{n,r})+\Theta(n)$ and $T_{n,r}\subseteq G$ for every $G\in\mathrm{EX}(n,\mathcal{F})$, determine whether or not we have $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$.
We answer this question in the negative for all $r\geq 3$ (see the proof of 1).
1.1 Notation and definitions
We write $K_{n}$ for the complete graph on $n$ vertices; $K_{V_{1},\ldots,V_{r}}$ for the complete multipartite graph with partite sets $V_{1},\ldots,V_{r}$; $T_{n,r}$ for the $r$-partite Turán graph on $n$ vertices; $P_{\ell}$ for a path on $\ell$ vertices; and $M_{n}$ for a matching on $n$ vertices (or almost matching if $n$ is odd). For a graph family $\mathcal{F}$, let $\chi(\mathcal{F})=\min\{\chi(F):F\in\mathcal{F}\}$. Given a graph $G$, we write $V(G)$ and $E(G)$ for its vertex set and edge set, respectively; given a set $S\subseteq V(G)$, we write $N_{G}(S):=\{v\in V(G):v\sim u\text{ for all }u\in S\}$, or just $N(S)$ when the underlying graph is understood; $G[U]$ for the induced subgraph on $U$, $E[U]$ for $E(G[U])$ and $e(U)$ for $|E(U)|$; $G[U,W]$ for the bipartite subgraph induced by $U$ and $W$; $A(G)$ for the adjacency matrix of $G$ and $\lambda(G)$ for the spectral radius of $A(G)$. For a vertex $v\in V(G)$ and a set $S\subseteq V(G)$, we write $N_{S}(v):=N(v)\cap S$ and $d_{S}(v)$ for $|N_{S}(v)|$; we write $G-v$ for the graph obtained by deleting the vertex $v$; and $d^{\prime}(v)$ for the number of walks of length two starting from $v$. For $n\in\mathbb{N}$, let $U(G,n)$ be a graph on $n$ vertices consisting of as many disjoint copies of $G$ as possible with isolated vertices in the remainder. We write $G\cup H$ and $G+H$ for the disjoint union and join, respectively, of $G$ and $H$ and $k\cdot G$ for the union of $k$ disjoint copies of $G$.
1.2 Main results
Our main result extends Theorem 1 in the case $r\geq 3$. Let $\alpha$ be the largest root of the quadratic
$f(x)=-\frac{1}{r-1}x^{2}+\left(2+\frac{5}{r-1}-\frac{4}{r}\right)x+\left(\frac{4}{r}-\frac{6}{r-1}\right),$
and let $c_{1}(r)=\frac{\alpha-1}{\alpha r},$ i.e.
$c_{1}(r)=\frac{1}{r}\left(1-\left[\frac{r-1}{2}\left(2+\frac{5}{r-1}-\frac{4}{r}+\sqrt{\left(2+\frac{5}{r-1}-\frac{4}{r}\right)^{2}-\frac{4}{r-1}\left(\frac{6}{r-1}-\frac{4}{r}\right)}\right)\right]^{-1}\right).$
Theorem 2.
Let $\mathcal{F}$ be a finite graph family satisfying $\chi(\mathcal{F})=r+1\geq 4$ and $\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,r})+Qn$ for large enough $n$, where $Q<c_{1}(r)$. Then we have $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$ for large enough $n$.
The appearance of the constant $c_{1}(r)$ is explained as follows. The ‘worst case scenario’ in the proof of Theorem 2 occurs when an edge extremal graph $H$ is obtained from $T_{n,r}$ by embedding paths $P_{k}$ in a large part along with one extra edge, while a spectral extremal graph $G$ is obtained from $T_{n,r}$ by embedding stars $K_{1,k-1}$ in a small part. (The $k$-vertex trees with smallest and largest spectral radius are $P_{k}$ and $K_{1,k-1}$, respectively, so intuitively they should also give the extremal contributions to the spectral radius of a larger graph in which they are embedded.) We will show that
$\lambda(H)-\lambda(G)=\left[2-\frac{k-5+6/k}{r-1}-\frac{4(k-1)}{kr}\right]\frac{1}{n}+O(n^{-2}).$
The first term inside the brackets accounts for the extra edge; the second term comes from the difference between embedding stars versus paths; and the last term comes from the difference between embedding the trees in a small part versus a large part. In order to avoid this potential counterexample to Theorem 2, we require that the bracketed expression be positive, which is equivalent to $k<\alpha$. The number of edges embedded in the large part of $H$ is $\frac{k-1}{kr}n+O(1)<c_{1}(r)n$. Regarding the other possibility $k>\alpha$, we can always construct an $\mathcal{F}$ which gives rise to this ‘worst case scenario,’ which proves that the constant $c_{1}(r)$ in Theorem 2 is sharp up to some integrality considerations:
Proposition 1.
Suppose that $r\geq 3$, $k\in\mathbb{N}$, and
$k>\frac{r-1}{2}\left(2+\frac{5}{r-1}-\frac{4}{r}+\sqrt{\left(2+\frac{5}{r-1}-\frac{4}{r}\right)^{2}-\frac{4}{r-1}\left(\frac{6}{r-1}-\frac{4}{r}\right)}\right).$
Then there is a finite family of graphs $\mathcal{F}$ such that $\chi(F)=r+1$, $\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,r})+\frac{k-1}{kr}n+O(1)$, and $\mathrm{SPEX}(n,\mathcal{F})\cap\mathrm{EX}(n,\mathcal{F})=\emptyset$ for an infinite sequence of $n$. Thus, $c(r)\leq\frac{k-1}{kr}.$
Corollary 1.
If $r\geq 3$ and
$\frac{r-1}{2}\left(2+\frac{5}{r-1}-\frac{4}{r}+\sqrt{\left(2+\frac{5}{r-1}-\frac{4}{r}\right)^{2}-\frac{4}{r-1}\left(\frac{6}{r-1}-\frac{4}{r}\right)}\right)$
is not an integer, then $c(r)=\frac{k-1}{kr}$, where $k$ is the ceiling of the expression above.
We illustrate 1 by giving some precise values of $c(r)$.
$r$ $3$ $4$ $5$ $6$ $7$ $8$ $9$ $10$
$c(r)$ $5/18$ $7/32$ $9/50$ $11/72$ $13/98$ $15/128$ $17/162$ $19/200$
Additionally, the asymptotics of $c(r)$ can be deduced from Theorem 2 and 1, namely $c(r)\sim 1/r$ as $r\to\infty$.
The assumption in Theorem 2 that $r\geq 3$ is required, as shown by the following result.
Proposition 2.
There exists a finite graph family $\mathcal{F}$ such that $\chi(F)=3$, $\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,2})+\frac{1}{3}n$, and $\mathrm{SPEX}(n,\mathcal{F})\cap\mathrm{EX}(n,\mathcal{F})=\emptyset$ for an infinite sequence of $n$.
We conjecture that a weaker result holds when $r=2$.
Conjecture 1.
Let $n$ be sufficiently large and let $\mathcal{F}$ be a finite graph family satisfying $\chi(F)=3$ and $\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,2})+Qn$, where $Q<\frac{1}{3}$. Then for $n$ large enough we have $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$.
1.3 Applications
Theorem 2 unifies several new and existing results. Taking $Q=\frac{1}{2r}$ in Theorem 2, one may obtain the case $r\geq 3$ of the previous results [wang2023conjecture] and [fang2024turan]. Using this, one can prove that $\mathrm{SPEX}(n,F)\subseteq\mathrm{EX}(n,F)$ for $n$ large enough when $F$ is a $(k,r)$-fan obtained by intersecting $k$ copies of $K_{r}$ in a single vertex (where $r\geq 4$) [cioaba2019spectral, desai2022spectral].
To obtain new applications of Theorem 2, we are interested in families $\mathcal{F}$ such that $e(T_{n,r})+\left\lfloor\frac{n}{2r}\right\rfloor\leq\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,r})+Qn$. One example is the graph $P_{2p+1}^{p}$ obtained by joining all vertices at distance at most $p$ in a path of length $12p+$ 1, where $p\geq 3$. Then Theorem 1.2 of [yuan2022extremal] gives that $\mathrm{EX}(n,\mathcal{F})$ contains a graph obtained from $T_{n,p}$ with a matching embedded in one part, hence Theorem 2 implies that $\mathrm{SPEX}(n,P_{2p+1}^{p})\subseteq\mathrm{EX}(n,P_{2p+1}^{p})$ for $n$ large enough.
We are not aware of any other existing extremal results in this category. However, with the goal of constructing graphs $F$ for which we can show $\mathrm{ex}(n,F)=e(T_{n,r})+\frac{n}{2r}+O(1)$ we began exploring specific small graphs. Other than the path power discussed above, the simplest such $F$ we found is described as follows. The graph $F_{1}$ consists of two paths on four vertices $bcde$, $fghi$, with additional edges $bf,bg,cf,ch,dg,di,eh,ei$ and additionally one dominating vertex $a$.
abcdefghi
Figure 1: The graph $F_{1}$.
Claim 1.
We have the following properties of $F_{1}$.
  • (i)
    $\chi(F_{1})=4$.
  • (ii)
    The graph obtained from $T_{n,3}$ by adding an edge to two different parts contains a copy of $F_{1}$.
  • (iii)
    The graph obtained from $T_{n,3}$ by embedding $P_{3}\cup P_{2}$ in part contains a copy of $F_{1}$.
  • (iv)
    The graph obtained from $T_{n,3}$ by embedding a matching in one part does not contain a copy of $F_{1}$.
Proof.
  • (i)
    One can check that $\chi(F_{1}-a)=3$ so $\chi(F_{1})=4$.
  • (ii)
    If the added edges are $u_{1}v_{2}$ and $v_{1}v_{2}$ then we find a copy of $F_{1}$ by identifying $bf$ with $u_{1}u_{2}$ and $ei$ with $v_{1}v_{2}$.
  • (iii)
    If the added paths are $u_{1}u_{2}u_{3}$ and $v_{1}v_{2}$ then we find a copy of $F_{1}$ by identifying $gbc$ with $u_{1}u_{2}u_{3}$ and $ei$ with $v_{1}v_{2}$.
  • (iv)
    Suppose $F_{1}$ appears in this graph. Assume without loss of generality that the matching is embedded in $V_{1}$, $a\in V_{3}$, and $b\in V_{1}$. If $f\in V_{1}$ then $c,g\in V_{2}$ so $d,h\in V_{1}$ and so $e,i\in V_{2}$, contradicting $e\sim i$. If $f\in V_{2}$ then $c,g\in V_{1}$ so $gbc$ is a $P_{3}$ in $V_{1}$, a contradiction.
Facts (i), (ii), and (iii) give that the assumptions of 1 below are satisfied, hence the conclusion of 1 and (iv) give that $\mathrm{EX}(n,F_{1})$ is obtained by embedding a matching in one part of $T_{n,3}$. This gives a value of $Q=1/2r$ in Theorem 2, and we conclude that $\mathrm{SPEX}(n,F_{1})=\mathrm{EX}(n,F_{1})$.
2 Proof of Theorem 2
Our overall strategy is as follows. In subsection 2.1 we prove 1, which provides general structural information about the graphs in $\mathrm{EX}(n,\mathcal{F})$. In subsection 2.2, we prove that the same general structural description also applies to graphs in $\mathrm{SPEX}(n,\mathcal{F})$. The difference in spectral radius of two graphs $H,G$ satisfying this structural description is equal to $2(e(H)-e(G))/n+O(n^{-2})$, plus an error term which depends on what trees are embedded in $H$ and $G$ and in which part of their $r$-partition. We show that this error term is at most $\left[-\frac{k-5+6/k}{r-1}-\frac{4(k-1)}{kr}\right]\frac{1}{n}+O(n^{-2})$. If $e(G)=e(H)$ then $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$, while otherwise $e(G)<e(H)$ and we have
$\lambda(H)-\lambda(G)\geq\left[2-\frac{k-5+6/k}{r-1}-\frac{4(k-1)}{kr}\right]\frac{1}{n}+O(n^{-2}).$
The assumption $Q<c_{1}(r)$ in Theorem 2 guarantees that the bracketed term is positive, giving a contradiction.
Definition 1.
Given a graph family $\mathcal{F}$ with $\chi(F)=r+1\geq 3$ and $\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,r})+Qn$ for large enough $n$ where $Q<1/r$, let $k=k(\mathcal{F})$ be the maximum integer such that for some tree $T_{k}$ of order $k$, for all $m$, $U(T_{k},m)+T_{m(r-1),r-1}$ is $\mathcal{F}$-free.
Since $Q<1/r$, $k$ is finite; more precisely, whenever $k|m$ we have $m\frac{k-1}{k}\leq Qmr$, hence $k\leq\lfloor 1/(1-Qr)\rfloor$.
2.1 Structure of edge extremal graphs
In this subsection we prove the following lemma. Note that if $\mathcal{F}$ is as in Theorem 2, then $\mathcal{F}$ satisfies the assumptions of 1. Our reason for using this lemma which is slightly more general than what is needed to prove Theorem 2 is that it also helps us quickly determine the extremal graphs for the forbidden subgraphs appearing in our applications and counterexamples.
Lemma 1.
Suppose $\mathcal{F}$ is a finite family of graphs, $\chi(\mathcal{F})=r+1\geq 3$, and for $n$ large enough, any graph obtained by adding more than $Qnr$ edges to $T_{nr,r}$ is not $\mathcal{F}$-free, where $Q<1/r$. Then for $n$ large enough, $\mathrm{EX}(n,\mathcal{F})$ contains a graph $G$ obtained from $T_{n,r}$ by embedding disjoint trees of some fixed order $k$ in one part, then making a bounded number of edge deletions and additions to the whole graph.
Let $k=k(\mathcal{F})$ as in 1.
Lemma 2.
For some large enough constant $m$, the following graphs are not $\mathcal{F}$-free:
  • $K_{1,m}+T_{m(r-1),r-1}$
  • $M_{m}+M_{m}+T_{m(r-2),r-2}$
  • $(T_{1}\cup\cdots\cup T_{m})+T_{m(r-1),r-1}$, for any trees $T_{i}$ of order $k+1$
  • $(C_{1}\cup\cdots\cup C_{m})+T_{m(r-1),r-1}$, for any cycles $C_{i}$ of order at most $k$
Proof.
Assume for a contradiction one of the graphs above is $\mathcal{F}$-free.
  • $K_{1,m}+T_{m(r-1),r-1}$ contains a graph obtained by adding $m-1$ edges to $T_{m(r-1),r-1}$, so $Q(m-1)r\geq m$ by assumption and $Q\geq\left(1-\frac{1}{m}\right)/r$, which contradicts $Q<1/r$ for large enough $m$.
  • $M_{m}+M_{m}+T_{m(r-2),(r-2)}$ is obtained by adding $m$ edges to $T_{mr,r}$, so $Qmr\geq m$, hence $Q\geq 1/r$, a contradiction.
  • By applying Cayley’s formula and the pigeonhole principle, $(T_{1}\cup\cdots\cup T_{m})+T_{m(r-1),r-1}$ contains a graph obtained by embedding $U(T,m/(k+1)^{k})$ in one part of $T_{mr/(k+1)^{k},r}$ for some tree $T$ of order $k+1$. If $m$ is large enough, this violates the definition of $k(\mathcal{F})$.
  • Let $m^{\prime}$ be the sum of the orders of the $C_{i}$, then $m^{\prime}\geq m$ and $(C_{1}\cup\cdots\cup C_{m})+T_{m^{\prime}(r-1),r-1}$ is obtained from $T_{m^{\prime}r,r}$ by adding $m^{\prime}$ edges to one part, which contradicts $Q<1/r$.
Assume that $G=G_{n}$ is any sequence of graphs in $\mathrm{EX}(n,\mathcal{F})$. Let $m$ be a large enough even constant satisfying 2. Let $M=\max\{|V(F)|:F\in\mathcal{F}\}$. Choose constants $\varepsilon,\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}$ satisfying the following inequalities (we do not attempt to avoid redundancy or optimize the constants).
(i) $M\sqrt{\varepsilon}<1/r$ (ii) $\sqrt{\varepsilon}+2\varepsilon_{1}<1/r$ (iii) $\sqrt{\varepsilon}+\varepsilon_{1}<\varepsilon_{2}$ (iv) $\sqrt{\varepsilon}<\varepsilon_{1}$ (v) $(2mr+1)\sqrt{\varepsilon}<\varepsilon_{1}$ (vi) $\varepsilon_{2}+2\sqrt{\varepsilon}<\varepsilon_{3}$ (vii) $\sqrt{\varepsilon}+m(k+1)\varepsilon_{3}<1/r$ (viii) $\sqrt{\varepsilon}+m(k+r)\varepsilon_{3}<1/r$ (ix) $\sqrt{\varepsilon}+M\varepsilon_{3}<1/r$ (x) $\sqrt{\varepsilon}+2mr\varepsilon_{3}<1/r$
This is possible, as seen by choosing them in the order $\varepsilon_{3},\varepsilon_{2},\varepsilon_{1},\varepsilon$.
The Erdős-Simonovits stability theorem gives the following lemma.
Lemma 3 (Erdős-Simonovits [erdos1967some],[erdos1966some],[simonovits1968method]).
If $n$ is large enough then $G$ can be obtained from $T_{n,r}$ by adding and deleting at most $\varepsilon n^{2}$ edges.
Let $V(G)=V=V_{1}\cup\cdots\cup V_{r}$ where $|V_{i}|\in\{\lfloor n/r\rfloor,\lceil n/r\rceil\}$, and $G$ is obtained from $K_{V_{1},\ldots,V_{r}}$ by adding a graph $A$ and removing a graph $B$ with $e(A)+e(B)\leq\varepsilon n^{2}$. Let $X=\{v\in V:d_{B}(x)\geq 2\sqrt{\varepsilon}n\}$, and note that we have
$2|X|\sqrt{\varepsilon}n\leq\sum_{v\in X}d_{B}(v)=2e(B)\leq 2\varepsilon n^{2}\Longrightarrow|X|\leq\sqrt{\varepsilon}n.$
For $v\in V_{i}$, we say that $v$ is $V$-ordinary if $N(v)=\bigcup_{j\neq i}V_{j}-X$.
Lemma 4.
If $v\in X$, then there exists some $i$ such that $d_{V_{i}}(v)\leq\varepsilon_{1}n$ and $d_{V_{j}}(v)\geq n/r-\varepsilon_{2}n$ for every $j\neq i$.
Proof.
Suppose not. Note than a $V$-ordinary vertex $u$ satisfies $d(u)\geq(1-1/r-\sqrt{\varepsilon})n+O(1)$.
Claim 2.
If $G^{\prime}$ is obtained from $G$ by deleting some vertex and adding a $V$-ordinary vertex $u\in V_{i}$, then $G^{\prime}$ is $\mathcal{F}$-free.
Proof.
Suppose $G^{\prime}$ contains some $F\in\mathcal{F}$. Then $u\in V(F)$, and $|N_{F}(u)|\leq M$, $N_{F}(u)\subseteq V-V_{i}-X$, so
$|N_{G}(N_{F}(u))|\geq(1/r-M\sqrt{\varepsilon})n+O(1)>0$
since $M\sqrt{\varepsilon}<1/r$, (i). Thus replacing $u$ by a vertex in $N_{G}(N_{F}(u))$ gives a copy of $F$ in $G$, a contradiction. ∎
If there are two different indices $i,i^{\prime}$ such that $d_{V_{i}}(v),d_{V_{i^{\prime}}}(v)\leq\varepsilon_{1}n$ then we have
$d(v)\leq(1-2/r+2\varepsilon_{1})n+O(1).$
If we delete $v$ and add a $V$-ordinary vertex we obtain
$e(G^{\prime})-e(G)=(1/r-\sqrt{\varepsilon}-2\varepsilon_{1})n+O(1)>0$
since $\sqrt{\varepsilon}+2\varepsilon_{1}<1/r$ (ii), which contradicts $G\in\mathrm{EX}(n,\mathcal{F})$ by 2. If there is a unique index $i$ such that $d_{V_{i}}(v)\leq\varepsilon_{1}n$ then by assumption there exists some $j\neq i$ such that $d_{V_{j}}(v)\leq n/r-\varepsilon_{2}n$. Thus
$d(v)\leq(1-1/r+\varepsilon_{1}-\varepsilon_{2})n+O(1)$
and if we delete $v$ and add an $V$-ordinary vertex we obtain
$e(G^{\prime})-e(G)=(\varepsilon_{2}-\varepsilon_{1}-\sqrt{\varepsilon})n+O(1)>0$
since $\sqrt{\varepsilon}+\varepsilon_{1}<\varepsilon_{2}$ (iii), which contradicts $G\in\mathrm{EX}(n,\mathcal{F})$ by 2. Finally, suppose there is no index $i$ for which $d_{V_{i}}(v)\leq\varepsilon_{1}n$, and assume without loss of generality $v\in V_{1}$. Now
$|N(v)\cap(V_{1}-X)|\geq(\varepsilon_{1}-\sqrt{\varepsilon})n+O(1)\geq m$
since $\sqrt{\varepsilon}<\varepsilon_{1}$ (iv), so there are some $m$ vertices $v_{11},\ldots,v_{1m}\in N(v)\cap(V_{1}-X)$. Having found $v_{11},\ldots,v_{1m},$ $\ldots,v_{i1},\ldots,v_{im}$, we have
$|N(\{v,v_{11},\ldots,v_{im}\})\cap V_{i+1}-X|\geq(\varepsilon_{1}-2im\sqrt{\varepsilon}-\sqrt{\varepsilon})n+O(1)\geq m$
since $(2mr+1)\sqrt{\varepsilon}<\varepsilon_{1}$ (v), so we can continue this process until we have selected $v_{11},\ldots,v_{rm}$. Then $G[\{v,v_{11},\ldots,v_{rm}\}]\supseteq K_{1,m}+T_{m(r-1),r-1}$, contradicting 2. ∎
By 4, we may define a new partition $V=W_{1}\sqcup\cdots\sqcup W_{r}$ as follows. If $v\in V_{i}-X$ then let $v\in W_{i}$. If $v\in X$ then let $v\in W_{i}$ where $i$ is chosen such that $d_{V_{i}}(v)\leq\varepsilon_{1}n$ and $d_{V_{j}}(v)\geq n/r-\varepsilon_{2}n$ for every $j\neq i$. Let $w_{i}:=|W_{i}|$. Note that $w_{i}\in(n/r-\sqrt{\varepsilon}n+O(1),n/r+\sqrt{\varepsilon}n+O(1))$. Thus for each $v\in W_{i}$, for each $j\neq i$, $d_{W_{j}}(v)\geq w_{j}-(\varepsilon_{2}+2\sqrt{\varepsilon})n+O(1)\geq w_{j}-\varepsilon_{3}n$, since $\varepsilon_{2}+2\sqrt{\varepsilon}<\varepsilon_{3}$, (vi).
Lemma 5.
Let $K$ be a graph of order at most $m(k+1)$. For $n$ large enough, if $G[W_{i}]\supseteq K$ for some $i$ then $G\supseteq K+T_{m(r-1),r-1}$.
Proof.
Assume without loss of generality $i=1$, then
$|N(V(K))\cap W_{2}|\geq w_{2}-|V(K)|\varepsilon_{3}n+O(1)\geq(1/r-\sqrt{\varepsilon}-m(k+1)\varepsilon_{3})n+O(1)\geq m$
since $\sqrt{\varepsilon}+m(k+1)\varepsilon_{3}<1/r$, (vii), so assume $v_{21},\ldots,v_{2m}\in N(V(K))\cap W_{2}$. Having found $v_{21},\ldots,v_{im}$, we have
$|N(V(K)\cup\{v_{21},\ldots,v_{im}\})\cap W_{i+1}|$ $\geq w_{i+1}-(|V(K)|+(i-1)m)\varepsilon_{3}n+O(1)$
$\geq(1/r-\sqrt{\varepsilon}-m(k+r)\varepsilon_{3})n+O(1)\geq m$
since $\sqrt{\varepsilon}+m(k+r)\varepsilon_{3}<1/r$, (viii). Thus, we can continue the process until we have found $v_{21},\ldots,v_{rm}$, and $G[V(K)\cup\{v_{21},\ldots,v_{rm}\}]\supseteq K+T_{m(r-1),r-1}$. ∎
Call a vertex $v\in W_{i}$ $W$-ordinary if $N(v)=V-W_{i}$.
Lemma 6.
Let $G^{\prime}$ be obtained from $G$ by deleting some vertex and adding an ordinary vertex $u\in W_{i}$. Then $G^{\prime}$ is $\mathcal{F}$-free.
Proof.
Suppose $G^{\prime}$ contains some $F\in\mathcal{F}$. Then $u\in V(F)$, $|N_{F}(u)|\leq M$ and $N_{F}(u)\subseteq V-W_{i}$. Hence
$|N_{G}(N_{F}(u))|\geq w_{i}-M\varepsilon_{3}n+O(1)\geq(1/r-\sqrt{\varepsilon}-M\varepsilon_{3})n+O(1)>0$
since $\sqrt{\varepsilon}+M\varepsilon_{3}<1/r$, (ix), so replacing $u$ by some vertex in $N_{G}(N_{F}(u))$ gives a copy of $F$ in $G$, a contradiction. ∎
Lemma 7.
For each $W_{i}$, for each $v\in W_{i}$, we have $d_{W_{i}}(v)\leq m$.
Proof.
This follows immediately from 5 and 2. ∎
Lemma 8.
For each $v\in W_{i}$, for each $j\neq i$, we have $d_{W_{j}}(v)\geq w_{j}-m$.
Proof.
If $d_{W_{j}}(v)\leq w_{j}-m-1$ then by 7 replacing $v$ with a $W$-ordinary vertex increases the number of edges, and by 6 and $G\in\mathrm{EX}(n,\mathcal{F}$) this is a contradiction. ∎
Lemma 9.
For some constant $C$, there is at most one $i$ such that $e(W_{i})\geq C$.
Proof.
Assume for a contradiction that $e(W_{1}),e(W_{2})$ are both arbitrarily large. By 7, $G[W_{1}],G[W_{2}]$ have bounded maximum degree, so this implies we can find a matching $M_{1}$ in $W_{1}$ of order $m$ and a matching $M_{2}$ in $W_{2}$ of order $2(m+1)m$. For each $v\in M_{1}$, by 8, there are at most $m$ edges in $M_{2}$ either of whose endpoints are not adjacent to $v$, which implies that there are some $(m+1)m-m^{2}=m$ edges in $M_{2}$ inducing a matching $M_{2}^{\prime}$ such that $G[M_{1},M_{2}^{\prime}]\simeq K_{m,m}$. Similarly to lemmas proved above, having found $v_{31},\ldots,v_{im}$, we have
$|N(M_{1}\cup M_{2}^{\prime}\cup\{v_{31},\ldots,v_{im}\})\cap W_{i+1}|$ $\geq w_{i+1}-(2m+(i-2)m)\varepsilon_{3}n+O(1)$
$\geq(1/r-\sqrt{\varepsilon}-2mr\varepsilon_{3})n+O(1)\geq m$
since $\sqrt{\varepsilon}+2mr\varepsilon_{3}<1/r$, (x). Thus, we find $M_{1},M_{2}^{\prime},v_{31},\ldots,v_{rm}$ whose union induces an $M_{m}+M_{m}+T_{(m-2)r,r}$, which contradicts 2. ∎
Assume without loss of generality that $W_{1}$ is the part for which $e(W_{1})\to\infty$ is possible, i.e. for all $i\neq 1$, we have $e(W_{i})=O(1)$. For each $i\neq 1$, we may write $W_{i}=W_{i}^{o}\sqcup W_{i}^{m}$, where $e(W_{i}^{o})=0$ and $|W_{i}^{m}|=O(1)$.
Lemma 10.
For each $i\neq 1$, all vertices in $W_{i}^{o}$ are $W$-ordinary.
Proof.
Let $v\in W_{i}^{o}$ not be ordinary. Since $d_{W_{i}}(v)=0$, replacing $v$ with an ordinary vertex increases the number of edges, which by 6 and $G\in\mathrm{EX}(n,\mathcal{F})$ is a contradiction. ∎
Lemma 11.
We can write $W_{1}=W_{1}^{o}\sqcup W_{1}^{m}$, where $G[W_{1}^{o}]$ is a disjoint union of trees of order $k$, each tree component of $G[W_{1}^{o}]$ occurs more than $M(m+1)$ times, and $|W_{1}^{m}|=O(1)$.
Proof.
We prove the result in a series of a claims.
Claim 3.
The number of components of $G[W_{1}]$ whose order is at least $k+1$ is at most $m$.
Proof.
If not, then we can find trees $T_{1},\ldots,T_{m}$ of order $k+1$ such that $T_{1}\cup\cdots\cup T_{m}\subseteq G[W_{1}]$. Then by 5, we have $(T_{1}\cup\cdots\cup T_{m})+T_{m(r-1),r-1}\subseteq G$, contradicting 2. ∎
Claim 4.
The graph $G[W_{1}]$ contains at most $m$ disjoint cycles of order at most $k$.
Proof.
If not, then let the $m$ cycles be called $C_{1},\ldots,C_{m}$. By 5, we have $(C_{1}\cup\cdots\cup C_{m})+T_{m(r-1),r-1}\subseteq G$, contradicting 2. ∎
Claim 5.
The graph $G[W_{1}]$ has at least $\sqrt{n}$ isomorphic tree components of order $k$.
Proof.
Suppose not. Then by 3, 4, and Cayley’s formula, all but $k^{k-2}\sqrt{n}+O(1)$ of the components of $G[W_{1}]$ are smaller trees. By the definition of $k(\mathcal{F})$, it is easy to check that the graph $G^{\prime}$ obtained by deleting $E(W_{i})$ for all $i$ and then embedding $U(T_{k},w_{1})$ in $W_{1}$ for some $k$-vertex tree $T_{k}$ is $\mathcal{F}$-free. Then we have
$e(G^{\prime})-e(G)\geq w_{1}\frac{k-1}{k}+O(1)-k^{k-2}\sqrt{n}-w_{1}\frac{k-2}{k-1}>0$
which contradicts $G\in\mathrm{EX}(n,\mathcal{F})$. ∎
Let $T^{\prime}$ be a tree occurring at least $\sqrt{n}$ times as guaranteed by 5.
Claim 6.
For any set $S\subseteq W_{1}$ with $|S|=k$, the graph $G^{\prime}$ obtained by deleting $E(S,W_{1})$ and adding a copy of $T^{\prime}$ on $S$ is $\mathcal{F}$-free.
Proof.
Assume for a contradiction that $G^{\prime}\supseteq F\in\mathcal{F}$. Then $S^{\prime}:=V(F)\cap S\neq\emptyset$. Note there are at least $\sqrt{n}/2$ copies of $T^{\prime}$ which do not intersect $F$ in any vertex. Now we have
$|N(N_{F}(S^{\prime}))\cap W_{1}|\geq w_{1}-Mm\geq w_{1}-\sqrt{n}/4$
using 8, so there is a copy of $T^{\prime}$ such that $V(T^{\prime})\cap V(F)=\emptyset$ and $V(T^{\prime})\subseteq N(N_{F}(S^{\prime}))$. Replacing $S^{\prime}$ with the corresponding vertices in $T^{\prime}$ gives a copy of $F$ in $G$, a contradiction. ∎
Claim 7.
If the tree components of $G[W_{1}]$ of order at most $k-1$ are $T_{1},\ldots,T_{j}$, then we have
$\sum_{i=1}^{j}|V(T_{i})|\leq(k-1)^{2}k.$
Proof.
If not, then for some $\ell\leq k-1$, there are $k$ tree components $T_{1},\ldots,T_{k}$ of order $\ell$. By 6, if $G^{\prime}$ is obtained from $G$ by replacing $T_{1}\cup\cdots\cup T_{k}$ with $\ell\cdot T^{\prime}$, then $G^{\prime}$ is $\mathcal{F}$-free. Also, $e(G^{\prime})>e(G)$ which contradicts $G\in\mathrm{EX}(n,\mathcal{F})$. ∎
Combining 3, 4, and 7 gives the result (we put any trees of order $k$ which occur at most $M(m+1)$ times in $W_{1}^{m}$). ∎
Lemma 12.
Every vertex in $W_{1}^{o}$ is $W$-ordinary.
Proof.
Assume for a contradiction that $v\in W_{1}^{o}$ and $N(v)\not\supseteq V-W_{1}$, so there is some $u\not\in W_{1}$ with $v\not\sim u$. Then the graph $G^{\prime}$ obtained by adding the edge $uv$ contains some $F\in\mathcal{F}$. Let the component of $G[W_{1}]$ containing $v$ be $T$. By 11, there are more than $M(m+1)$ disjoint copies of $T$ in $G[W_{1}]$, and more than $Mm$ of these do not intersect $F$ in any vertex. Then from 8 we have
$|N(N_{F}(v))\cap W_{1}|\geq w_{1}-Mm$
so some copy of $T$ is contained in $N(N_{F}(v))$. This implies $F\subseteq G$, a contradiction. ∎
At this point we know that $G$ can be obtained from $K_{W_{1},\ldots,W_{r}}$ by embedding disjoint trees of order $k$ in $W_{1}$, then deleting and adding $O(1)$ edges.
Lemma 13.
Suppose $H$ is any graph obtained from $K_{W_{1},\ldots,W_{r}}$ by embedding disjoint trees in $W_{1}$, then deleting and adding $O(1)$ edges. Let $W_{i}^{o},W_{i}^{m}$ be defined as above, and assume that $|W_{i}^{o}|\to\infty$, $|W_{i}^{m}|=O(1)$ for each $i$. For any $i,j$, let $H_{ij}$ be defined by deleting a vertex $u$ in $W_{i}^{o}$ such that $d_{W_{i}}(u)\leq 1$ and adding a vertex $v$ whose neighborhood is $V-W_{j}-u$. If $H$ is $\mathcal{F}$-free then so is $H_{ij}$. Moreover, $e(H_{ij})-e(G)=|W_{i}|-|W_{j}|-1$ if $i\neq 1$ and $e(H_{ij})-e(G)\geq|W_{i}|-|W_{j}|-2$ if $i=1$.
Proof.
Suppose $H_{ij}$ has a copy of some $F\in\mathcal{F}$. If $H$ is $\mathcal{F}$-free, we must have that $v\in V(F)$. Since $|W_{j}|\to\infty$ and $|V(F)|\leq M$, there is some $v^{\prime}\in W_{j}^{o}-V(F)$, and $N_{H}(v^{\prime})\supseteq N_{H_{ij}}(v)$. Thus, replacing $v$ by $v^{\prime}$ gives a copy of $F$ in $H$, a contradiction. The second claim is obvious. ∎
Below we refer to $H_{ij}$ for various graphs $H$ without restating the full setup when it is clear from context. We will make statements such as “$H$ satisfies $|w_{i}-w_{j}|\leq a$ for all $i,j$,” where the $w_{i}$ are assumed to be the values of $|W_{i}|$ which satisfy the assumptions of 13.
Lemma 14.
For any $i,j\in\{2,\ldots,r\}$, we have $|w_{i}-w_{j}|\leq 1$.
Proof.
If not, then there is some $i,j\geq 2$ such that $w_{i}\geq w_{j}+2$. By 13, $G_{ij}$ is $F$-free and $e(G_{ij})-e(G)=w_{i}-w_{j}-1\geq 1$, which is a contradiction. ∎
Lemma 15.
For any $i\in\{2,\ldots,r\}$, we have $w_{1}-2\leq w_{i}\leq w_{1}+1$.
Proof.
If $w_{1}-3\geq w_{i}$ then by 13 $G_{1i}$ is $\mathcal{F}$-free and $e(G_{1i})-e(G)\geq w_{1}-w_{i}-2\geq 1$, a contradiction. If $w_{i}\geq w_{1}+2$ then we have that $G_{i1}$ is $\mathcal{F}$-free and $e(G_{i1})-e(G)=w_{i}-w_{1}-1\geq 1$, a contradiction. ∎
Proof of 1.
By the above, we are done unless $w_{1}=w_{i}+2$ for some $i$. Then by 13, $G_{1i}$ is $\mathcal{F}$-free and $e(G_{1i})-e(G)\geq 0$, i.e. $G_{1i}\in\mathrm{EX}(n,\mathcal{F})$. One can check that $G_{1i}$ satisfies $|w_{i}-w_{j}|\leq 1$ for all $i,j$. ∎
2.2 Structure of spectral extremal graphs
We now prove Theorem 2. Let $\mathcal{F}$ be as in Theorem 2 and let $k=k(\mathcal{F})$ as in 1. If $k=1$, then the conclusion of Theorem 2 follows from Theorem 1, so we will assume that $k\geq 2$. Similarly to the previous subsection, we have the following lemma.
Lemma 16.
For some large enough even constant $m$, the following graphs are not $\mathcal{F}$-free.
  • $K_{1,m}+T_{m(r-1),r-1}$
  • $M_{m}+M_{m}+T_{m(r-2),r-2}$
  • $(T_{1}\cup\cdots\cup T_{m})+T_{m(r-1),r-1}$, for any trees $T_{i}$ of order $k+1$
  • $(C_{1}\cup\cdots\cup C_{m})+T_{m(r-1),r-1}$, for any cycles $C_{i}$ of order at most $k$
Let $m$ be a large enough even constant satisfying 16. Let $M=\max\{|V(F)|:F\in\mathcal{F}\}$. Suppose $G\in\mathrm{SPEX}(n,\mathcal{F})$, where $n$ is large enough. Choose positive constants $\varepsilon,\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}$,$\varepsilon_{4},\varepsilon_{5},\varepsilon_{6}$ satisfying the following inequalities (we do not attempt to avoid redundancy or optimize the constants).
(i) $2(r+1)\sqrt{\varepsilon}<\min\{\varepsilon_{1},\varepsilon_{3}\}$ (ii) $2\varepsilon_{1}+\sqrt{\varepsilon}<1/r$ (iii) $2(r+1)^{2}(\sqrt{\varepsilon}+\varepsilon_{1})<\varepsilon_{2}$ (iv) $M\sqrt{\varepsilon}<1/r$ (v) $\sqrt{\varepsilon}+\varepsilon_{2}+2\varepsilon_{3}<1/r$ (vi) $\sqrt{\varepsilon}+\varepsilon_{2}+\varepsilon_{3}<\varepsilon_{4}$ (vii) $(2mr+1)\sqrt{\varepsilon}<\varepsilon_{3}$ (viii) $\max\{\varepsilon_{4}+\sqrt{\varepsilon},3\sqrt{\varepsilon}\}<\varepsilon_{5}$ (ix) $\sqrt{\varepsilon}+m(k+1)\varepsilon_{5}<1/r$ (x) $\sqrt{\varepsilon}+m(k+r)\varepsilon_{5}<1/r$ (xi) $\sqrt{\varepsilon}+M\varepsilon_{5}<1/r$ (xii) $\frac{r(\varepsilon_{1}+(r-1)\varepsilon_{5})}{r-1}<\varepsilon_{6}$ (xiii) $\frac{r}{r-1}(\varepsilon_{5}+\varepsilon_{1}+3\sqrt{\varepsilon}+\varepsilon_{2}/r+(r-1)\varepsilon_{5})<\varepsilon_{6}$ (xiv) $(m+1)\varepsilon_{6}<1$ (xv) $\sqrt{\varepsilon}+2mr\varepsilon_{5}<1/r$ (xvi) $2\varepsilon_{6}<1-\frac{k(k-2)}{(k-1)^{2}}$ (xvii) $2\varepsilon_{6}<1-\frac{k(k-2)}{(k-1)^{2}}$.
This is possible, as seen by choosing them in the order $\varepsilon_{6},\varepsilon_{5},\varepsilon_{4},\varepsilon_{3},\varepsilon_{2},\varepsilon_{1},\varepsilon$.
We will use the spectral version of the stability theorem.
Lemma 17 (Desai-Kang-Li-Ni-Tait-Wang [desai2022spectral]).
If $n$ is large enough then $G$ can be obtained from $T_{n,r}$ by adding and deleting at most $\varepsilon n^{2}$ edges.
Therefore, let $V(G)=V_{1}\sqcup\cdots\sqcup V_{r}$, where $|V_{i}|\in\{\lfloor n/r\rfloor,\lceil n/r\rceil\}$ and $G$ is obtained from $K_{V_{1},\ldots,V_{r}}$ by adding a graph $A$ and removing a graph $B$ with $e(A)+e(B)\leq\varepsilon n^{2}$. Let $X=\{v\in V:d_{B}(x)\geq 2\sqrt{\varepsilon}n\}$ and note that we immediately have $|X|\leq\sqrt{\varepsilon}n$.
The proof of the following lemma follows the lines of Lemma 3.1 of [wang2023conjecture].
Lemma 18.
The graph $G$ is connected.
Proof.
Suppose $G$ is not connected. Then $\lambda(G)=\lambda(G_{1})$, where $G_{1}$ is a largest component of $G$. Let $G^{\prime}$ be obtained from $G_{1}$ by choosing any vertex $v\in V(G_{1})$ and adding a vertex $u$ pendant to $v$, along with $n-|V(G_{1})|-1$ isolated vertices. Then $\lambda(G_{1})>\lambda(G)\geq\lambda(T_{n,r})\to\infty$ and $G_{1}$ contains a copy of some $F\in\mathcal{F}$. On the other hand, we claim that $d_{G_{1}}(v)\leq M$; otherwise we could replace $u$ with some other vertex to show that $G\supseteq F$. Since $v$ was arbitrary, we find that $\Delta(G_{1})\leq M$ and consequently $\lambda(G_{1})\leq M$, a contradiction. ∎
Let $x$ be the principal eigenvector of $G$, normalized so that its maximum entry is $1$. Let $1=x_{z}=\max\{x_{v}:v\in V\}$. By 18, the Perron-Frobenius theorem implies that all entries of $x$ are positive.
Lemma 19.
There exists a unique $i_{1}$ such that $d_{V_{i_{1}}}(z)\leq\varepsilon_{1}n$ and for any $j\neq i_{1}$, we have $d_{V_{j}}(z)\geq(1/r-\varepsilon_{1})n+O(1)$.
Proof.
First, note that $d(z)\geq\lambda x_{z}=\lambda\geq\left(1-\frac{1}{r}\right)n+O(1)$. We claim there is a unique $i\in[r]$ such that $d_{V_{i}-X}(z)\leq(\varepsilon_{1}-\sqrt{\varepsilon})n$ and for any $j\neq i$, $d_{V_{j}}(z)\geq\frac{1}{r}n-\varepsilon_{1}n$. Suppose there is no $i$ such that $d_{V_{i}-X}(z)\leq(\varepsilon_{1}-\sqrt{\varepsilon})n$. By (i), there are some $m$ neighbors of $z$ in $V_{1}-X$, say $v_{11},\ldots,v_{1m}$. Having found $v_{11},\ldots,v_{im}$, we have
$|N(\{v,v_{11},\ldots,v_{im}\})\cap(V_{i+1}-X))|\geq(\varepsilon_{1}-\sqrt{\varepsilon}-2im\sqrt{\varepsilon}-\sqrt{\varepsilon})n\geq m$
since $2(r+1)\sqrt{\varepsilon}<\varepsilon_{1}$, (i). Thus, we eventually find a copy of $K_{1,m}+T_{m(r-1),r-1}$ in $G$, contradicting 16. So some such $i_{1}$ exists. To prove the uniqueness, suppose there are $i\neq j$ such that $d_{V_{i}-X}(z),d_{V_{j}-X}(z)\leq\varepsilon_{1}n$; then we have $d(z)\leq\left(1-\frac{2}{r}+2\varepsilon_{1}\right)n+O(1)$ which is a contradiction by (ii). Hence, there is a unique $i$ such that $d_{V_{i}-X}(z)\leq\varepsilon_{1}n$. Finally, since $d(z)\geq\left(1-\frac{1}{r}\right)n+O(1)$, we have that for any $j\neq i$, we have $d_{V_{j}}(z)\geq(1/r-\varepsilon_{1})n+O(1)$. ∎
Note that $\lambda\geq\lambda(T_{n,r})=\frac{r-1}{r}n+O(1)$.
Lemma 20.
For any $v\in V-X$, we have $x_{v}\geq 1-\varepsilon_{2}$.
Proof.
First suppose $v\in V_{i_{1}}$. Since $v\not\in X$, we have
$|N(z)-N(v)|\leq\varepsilon_{1}n+2(r-1)\sqrt{\varepsilon}n.$
and so
$\lambda x_{v}$ $\geq\lambda x_{z}-\sum_{u\in N(z)-N(v)}x_{u}$
$\geq\lambda x_{z}-(\varepsilon_{1}+2(r-1)\sqrt{\varepsilon})n$
$x_{v}$ $\geq 1-(\varepsilon_{1}+2(r-1)\sqrt{\varepsilon})\frac{n}{\lambda}$
$\geq 1-\frac{2r-1}{r-1}\varepsilon_{1}$
$\geq 1-\frac{2r^{2}\varepsilon_{1}}{r-1}+o(1)\geq 1-2r^{2}\varepsilon_{1}.$
Now if $v\not\in V_{i_{1}}$, then we have
$|N(z)-N(v)|\leq(1/r+2r\sqrt{\varepsilon})n+O(1)$
and
$|(N(v)-N(z))\cap V_{i_{1}}|\geq(1/r-2\sqrt{\varepsilon}-\varepsilon_{1})+O(1).$
Thus, applying the eigenweight bound above for vertices in $V_{i_{1}}$, we have
$\lambda x_{v}$ $\geq\lambda x_{z}-\sum_{u\in N(z)-N(v)}x_{u}+\sum_{u\in N(v)-N(z)}x_{u}$
$\geq\lambda-\frac{n}{r}-r2\sqrt{\varepsilon}n+\left(\frac{n}{r}-(2\sqrt{\varepsilon}+\varepsilon_{1})n\right)(1-2r^{2}\varepsilon_{1})+O(1)$
$\geq\lambda-\left(2(r+1)\sqrt{\varepsilon}+\varepsilon_{1}+2r\varepsilon_{1}\right)n+O(1)$
$\geq\lambda-2(r+1)(\sqrt{\varepsilon}+\varepsilon_{1})n+O(1)$
$x_{v}$ $\geq 1-2(r+1)(\sqrt{\varepsilon}+\varepsilon_{1})\frac{n}{\lambda}+o(1)\geq 1-\frac{2r(r+1)(\sqrt{\varepsilon}+\varepsilon_{1})}{r-1}+o(1)$
$\geq 1-2(r+1)^{2}(\sqrt{\varepsilon}+\varepsilon_{1}).$
The claim now follows from $2(r+1)^{2}(\sqrt{\varepsilon}+\varepsilon_{1})<\varepsilon_{2}$, (iii). ∎
The next section of the proof follows the lines of the argument in subsection 2.1. Define a $V$-ordinary vertex $v\in V_{i}$ to be one such that $N(v)=\bigcup_{j\neq i}V_{j}-X$.
Lemma 21.
Let $v\in X$. Then there exists some $i$ such that $d_{V_{i}}(v)\leq\varepsilon_{3}n$ and $d_{V_{j}}(v)\geq n/r-\varepsilon_{4}n$ for every $j\neq i$.
Proof.
Suppose not. Note that an ordinary vertex satisfies $d(u)\geq(1-1/r-\sqrt{\varepsilon})n+O(1)$.
Claim 8.
If $G^{\prime}$ is obtained from $G$ by deleting some vertex and adding a $V$-ordinary vertex $u\in V_{i}$, then $G^{\prime}$ is $\mathcal{F}$-free.
Proof.
We copy the proof of 2, using $M\sqrt{\varepsilon}<1/r$, (iv). ∎
If there are two different indices $i,i^{\prime}$ such that $d_{V_{i}}(v),d_{V_{i^{\prime}}}(v)\leq\varepsilon_{3}n$, then we have
$d(v)\leq(1-2/r+2\varepsilon_{3})n+O(1).$
By replacing $v$ with an ordinary vertex, we obtain that
$\frac{1}{2x_{v}}(x^{T}A(G^{\prime})x-x^{T}A(G)x)$ $\geq(1-1/r-\sqrt{\varepsilon})(1-\varepsilon_{2})n-(1-2/r+2\varepsilon_{3})n+O(1)$
$\geq(1/r-\sqrt{\varepsilon}-\varepsilon_{2}-2\varepsilon_{3})n+O(1)>0$
since $\sqrt{\varepsilon}+\varepsilon_{2}+2\varepsilon_{3}<1/r$, (v), which by 8 contradicts $G\in\mathrm{SPEX}(n,\mathcal{F})$. If there is a unique $i$ such that $d_{V_{i}}(v)\leq\varepsilon_{3}n$, then by assumption there is some $j$ for which $d_{V_{j}}(v)\leq n/r-\varepsilon_{4}n$. Thus
$d(v)\leq(1-1/r+\varepsilon_{3}-\varepsilon_{4})n+O(1).$
Then replacing $v$ by an ordinary vertex gives
$\frac{1}{2x_{v}}(x^{T}A(G^{\prime})x-x^{T}A(G)x)$ $\geq(1-1/r-\sqrt{\varepsilon})(1-\varepsilon_{2})n-(1-1/r+\varepsilon_{3}-\varepsilon_{4})n+O(1)$
$\geq(-\sqrt{\varepsilon}-\varepsilon_{2}-\varepsilon_{3}+\varepsilon_{4})n+O(1)>0$
since $\sqrt{\varepsilon}+\varepsilon_{2}+\varepsilon_{3}<\varepsilon_{4}$, (vi), which by 8 contradicts $G\in\mathrm{SPEX}(n,\mathcal{F})$. Finally, suppose there is no index $i$ for which $d_{V_{i}}(v)\leq\varepsilon_{3}n$. Then using $(2mr+1)\sqrt{\varepsilon}<\varepsilon_{3}$, (vii) and a similar argument to 4, we find a $K_{1,m}+T_{m(r-1),r-1}$, which contradicts 16. ∎
By 21, we may define a new partition $V=W_{1}\sqcup\cdots\sqcup W_{r}$ as follows. For $v\in V_{i}-X$, let $v\in W_{i}$. If $v\in X$, then let $v\in W_{i}$ where $i$ is chosen so that $d_{V_{i}}(v)\leq\varepsilon_{3}n$ and $d_{V_{j}}(v)\geq n/r-\varepsilon_{4}n$ for each $j\neq i$. Let $w_{i}:=|W_{i}|$. Note that $w_{i}\in(n/r-\sqrt{\varepsilon}n+O(1),n/r+\sqrt{\varepsilon}n+O(1))$. Thus for each $v\in W_{i}$, for each $j\neq i$, either $d_{W_{j}}(v)\geq w_{j}-(\varepsilon_{4}+\sqrt{\varepsilon})n+O(1)\geq w_{j}-\varepsilon_{5}n$ or $d_{W_{j}}(v)\geq w_{j}-(2\sqrt{\varepsilon}+\sqrt{\varepsilon})n+O(1)\geq w_{j}-\varepsilon_{5}n$, since $\max\{\varepsilon_{4}+\sqrt{\varepsilon},3\sqrt{\varepsilon}\}<\varepsilon_{5}$, (viii). Also, call a vertex $v\in W_{i}$ $W$-ordinary if $N(v)=V-W_{i}$.
Lemma 22.
Let $K$ be a fixed graph of order at most $m(k+1)$. For $n$ large enough, if $G[W_{i}]\supseteq K$ for some $i$ then $G\supseteq K+T_{m(r-1),r-1}$.
Proof.
We copy the proof of 5, using the inequalities $\sqrt{\varepsilon}+m(k+1)\varepsilon_{5}<1/r$, (ix), and $\sqrt{\varepsilon}+m(k+r)\varepsilon_{5}<1/r$, (x). ∎
Lemma 23.
Let $G^{\prime}$ be obtained from $G$ by deleting some vertex and adding a $W$-ordinary vertex $u\in V_{i}$. Then $G^{\prime}$ is $\mathcal{F}$-free.
Proof.
We copy the proof of 6, using the inequality $\sqrt{\varepsilon}+M\varepsilon_{5}<1/r$, (xi). ∎
Lemma 24.
For each $W_{i}$, for each $v\in W_{i}$, we have $d_{W_{i}}(v)\leq m$.
Proof.
This follows immediately from 22 and 16. ∎
Lemma 25.
For every $v\in V$, we have $x_{v}\geq 1-\varepsilon_{6}$.
Proof.
Let $z\in W_{i_{1}}$. If $v\in W_{i_{1}}$, then we have
$|N(z)-N(v)|\leq(\varepsilon_{1}+(r-1)\varepsilon_{5})n$
so
$\lambda x_{v}$ $\geq\lambda x_{z}-\sum_{u\in N(z)-N(v)}x_{u}$
$\geq\lambda-(\varepsilon_{1}+(r-1)\varepsilon_{5})n$
$x_{v}$ $\geq 1-\frac{(\varepsilon_{1}+(r-1)\varepsilon_{5})n}{\lambda}$
$\geq 1-\frac{r(\varepsilon_{1}+(r-1)\varepsilon_{5})}{r-1}+o(1).$
If $v\in W_{j}\neq W_{i_{1}}$, then we have
$|N(z)-N(v)|\leq w_{j}+(r-1)\varepsilon_{5}n\leq(1/r+\sqrt{\varepsilon}+(r-1)\varepsilon_{5})n+O(1)$
and
$|(N(v)-N(z))-X|\geq w_{i_{1}}-\varepsilon_{5}n-(\varepsilon_{1}+\sqrt{\varepsilon})n-\sqrt{\varepsilon}n+O(1)\\ \geq(1/r-\varepsilon_{5}-\varepsilon_{1}-2\sqrt{\varepsilon})n+O(1).$
Thus using 20 we have
$\lambda x_{v}$ $=\lambda x_{z}+\sum_{u\in N(v)-N(z)}x_{u}-\sum_{u\in N(z)-N(v)}x_{u}$
$\geq\lambda+(1/r-\varepsilon_{5}-\varepsilon_{1}-2\sqrt{\varepsilon})(1-\varepsilon_{2})n-(1/r+\sqrt{\varepsilon}+(r-1)\varepsilon_{5})n+O(1)$
$\geq\lambda-(\varepsilon_{5}+\varepsilon_{1}+3\sqrt{\varepsilon}+\varepsilon_{2}/r+(r-1)\varepsilon_{5})n+O(1)$
$x_{v}$ $\geq 1-\frac{r}{r-1}(\varepsilon_{5}+\varepsilon_{1}+3\sqrt{\varepsilon}+\varepsilon_{2}/r+(r-1)\varepsilon_{5})n+o(1)$
and the conclusion follows from $\frac{r(\varepsilon_{1}+(r-1)\varepsilon_{5})}{r-1}<\varepsilon_{6}$, (xii) and $\frac{r}{r-1}(\varepsilon_{5}+\varepsilon_{1}+3\sqrt{\varepsilon}+\varepsilon_{2}/r+(r-1)\varepsilon_{5})<\varepsilon_{6}$, (xiii). ∎
Lemma 26.
For each $v\in W_{i}$, for each $j\neq i$, we have $d_{W_{j}}(v)\geq w_{j}-m$.
Proof.
Suppose not. By 23, the graph $G^{\prime}$ obtained by replacing $v$ with a $W$-ordinary vertex is $\mathcal{F}$-free, and by 24 and 25 we have
$\frac{1}{2x_{v}}(x^{T}A(G^{\prime})x-x^{T}A(G)x)$ $\geq(m+1)(1-\varepsilon_{6})-m=1-(m+1)\varepsilon_{6}>0$
since $(m+1)\varepsilon_{6}<1$, (xiv), a contradiction. ∎
Lemma 27.
For some constant $C$, there is at most one $i$ such that $e(W_{i})\geq C$.
Proof.
We copy the proof of 9, using $\sqrt{\varepsilon}+2mr\varepsilon_{5}<1/r$ (xv). ∎
Assume without loss of generality that $W_{1}$ is the part for which $e(W_{i})\geq C$ is possible, so that $e(W_{i})=O(1)$ for all $i\neq 1$. Observe that for each $i\neq 1$, we may write $W_{i}=W_{i}^{o}\sqcup W_{i}^{m}$, where $e(W_{i}^{o})=0$ and $|W_{i}^{m}|=O(1)$. We will also define $W^{o}=\bigcup_{i=1}^{r}W_{i}^{o}$ and $W^{m}=\bigcup_{i=1}^{r}W_{i}^{m}$.
Lemma 28.
For each $i\neq 1$, for each $v\in W_{i}^{o}$, $N(v)=V-W_{i}$.
Proof.
We copy the proof of 10, using the fact that adding an edge to a connected graph increases its spectral radius. ∎
Lemma 29.
Suppose $G^{\prime}$ can be obtained from $G$ by deleting $a$ edges and adding $b$ edges, where $a/b<1-2\varepsilon_{6}$. Then $\lambda(G^{\prime})>\lambda(G)$.
Proof.
By 25, we have
$\frac{1}{2}(x^{T}A(G^{\prime})x-x^{T}A(G)x)\geq b(1-\varepsilon_{6})^{2}-a>b(1-2\varepsilon_{6})-a>0.$
Lemma 30.
We can write $W_{1}=W_{1}^{o}\sqcup W_{1}^{m}$, where $W_{1}^{o}$ is a union of trees of fixed order $k$, each tree component of $G[W_{1}^{o}]$ occurs more than $M(m+1)$ times, and $|W_{1}^{m}|=O(1)$.
Proof.
We prove the result in a series of claims.
Claim 9.
The number of components of $G[W_{1}]$ whose order is at least $k+1$ is at most $m$.
Proof.
We copy the proof of 3. ∎
Claim 10.
The graph $G[W_{1}]$ has at most $m$ disjoint cycles of order at most $k$.
Proof.
We copy the proof of 4. ∎
Claim 11.
The graph $G[W_{1}]$ has at least $\sqrt{n}$ tree components of order $k$.
Proof.
Similarly to 5, we find an $\mathcal{F}$-free graph $G^{\prime}$ obtained from $G$ by adding $b=w_{1}\frac{k-1}{k}+o(n)$ edges and deleting $a=w_{1}\frac{k-2}{k-1}+o(n)$ edges. Then
$\frac{a}{b}\leq\frac{k(k-2)}{(k-1)^{2}}+o(1)<1-2\varepsilon_{6}$
since $2\varepsilon_{6}<1-\frac{k(k-2)}{(k-1)^{2}}$, (xvi), so 29 gives $\lambda(G^{\prime})>\lambda(G)$, a contradiction. ∎
Let $T^{\prime}$ be a tree occurring at least $\sqrt{n}$ times as guaranteed by 11.
Claim 12.
For any set $S\subseteq W_{1}$ with $|S|=k$, the graph $G^{\prime}$ obtained by deleting $E(S)$ and adding a copy of $T^{\prime}$ on $S$ is $\mathcal{F}$-free.
Proof.
We copy the proof of 6. ∎
Claim 13.
If the tree components of $G[W_{1}]$ of order at most $k-1$ are $T_{1},\ldots,T_{a}$, then we have
$\sum_{i=1}^{a}|V(T_{i})|\leq(k-1)^{2}k.$
Proof.
As in 7, we obtain an $\mathcal{F}$-free graph $G^{\prime}$ obtained from $G$ by deleting $a=k(\ell-1)$ edges and adding $b=(k-1)\ell$ edges, where $\ell\leq k-1$. Then we have
$\frac{a}{b}\leq\frac{k(\ell-1)}{(k-1)\ell}\leq\frac{k(k-2)}{(k-1)^{2}}<1-2\varepsilon_{6}$
since $2\varepsilon_{6}<1-\frac{k(k-2)}{(k-1)^{2}}$, (xvii). Hence, 29 gives $\lambda(G^{\prime})>\lambda(G)$, contradicting $G\in\mathrm{SPEX}(n,\mathcal{F})$. ∎
Combining 9, 10, and 13, we obtain the result (we put any trees of order $k$ which occur at most $M(m+1)$ times in $W_{1}^{m}$). ∎
Lemma 31.
For each $v\in W_{1}^{o}$, $N(v)\supseteq V-W_{i}$.
Proof.
We copy the proof of 12, using the fact that adding an edge to a connected graph increases the spectral radius. ∎
Now we know that $G$ can be obtained from $K_{W_{1},\ldots,W_{r}}$ by embedding trees of order $k$ in $W_{1}$, then adding and deleting $O(1)$ edges. Now we determine the values of the $w_{i}$. Since all vertices in $W_{i}^{o}$ ($i\neq 1$) have the same neighborhood, they have the same eigenweight, so let $x_{i}$ be the common eigenweight of vertices in $W_{i}^{o}$. Let $x_{1}=\frac{1}{\lambda}\sum_{v\not\in W_{1}}x_{v}$. Observe that $\lambda=\Theta(n)$, a fact we will use frequently below.
Lemma 32.
For each $i\neq 1$, $v\in W_{i}$, we have $x_{v}=x_{i}+O(n^{-1})$.
Proof.
Let $u\in W_{i}^{o}$, and observe that $N(v)=N(u)\cup A^{\prime}-B^{\prime}$, where $a:=|A^{\prime}|=O(1)$ and $b:=|B^{\prime}|=O(1)$. Thus,
$\lambda x_{v}$ $=\lambda x_{u}+\sum_{w\in A^{\prime}}x_{w}-\sum_{w\in B^{\prime}}x_{b}$
$|x_{v}-x_{i}|$ $=\frac{1}{\lambda}\left|\sum_{w\in A^{\prime}}x_{w}+\sum_{w\in B}x_{w}\right|\leq\frac{1}{\lambda}(a+b)=O(n^{-1}).$
Lemma 33.
For any $1<i<j$, we have
$x_{j}=\frac{\lambda+w_{i}}{\lambda+w_{j}}x_{i}+O(n^{-2}).$
Proof.
We have
$\lambda x_{j}$ $=\lambda x_{i}+x_{i}(w_{i}-|W_{i}^{m}|)+\sum_{w\in W_{i}^{m}}x_{w}-x_{j}(w_{j}-|W_{j}^{m}|)-\sum_{w\in W_{j}^{m}}x_{w}$
$=\lambda x_{i}+w_{i}x_{i}+|W_{i}^{m}|O(n^{-1})-w_{j}x_{j}-|W_{j}^{m}|O(n^{-1})$
$=\lambda x_{i}+w_{i}x_{i}-w_{j}x_{j}+O(n^{-1}).$
Solving for $x_{j}$ gives the result. ∎
Lemma 34.
For $v\in W_{1}$, we have $x_{v}=x_{1}+O(n^{-1})$.
Proof.
We follow the proof of 32. ∎
Lemma 35.
For $v\in W_{1}$, we have $x_{v}=x_{1}+\frac{d_{W_{1}}(v)}{\lambda}x_{1}+O(n^{-2})$.
Proof.
We have
$\lambda x_{v}$ $=\lambda x_{1}+\sum_{u\in N_{W_{1}}(v)}x_{u}$
$=\lambda x_{1}+d_{W_{1}}(v)(x_{1}+O(n^{-1}))$
and since $\Delta(G[W_{1}])$ is bounded, solving for $x_{v}$ gives the result. ∎
The spectral version of 13 gives the following fact: for any $u\in W_{i}^{o}$, for any $j\neq i$, we have $\lambda(G_{ij})\leq\lambda(G)$ and so with $u$ being the transferred vertex, we have
$0\geq\frac{1}{2x_{u}}(x^{T}A(G_{ij})x-x^{T}A(G)x)=\sum_{v\in W_{i}-\{u\}-(N(u)\cap W_{i})}x_{v}-\sum_{v\in W_{j}}x_{v}$ (1)
For each $i$ write $X_{i}=\sum_{u\in W_{i}}x_{u}$. Note that for $i\neq 1$, $X_{i}=w_{i}x_{i}+O(n^{-1})$ and when $i=1$ we have
$X_{1}=\sum_{u\in W_{1}}(x_{1}+\frac{d_{W_{1}}(u)}{\lambda}x_{1}+O(n^{-2}))=w_{1}x_{1}+\frac{2e(W_{1})}{\lambda}x_{1}+O(n^{-1})=w_{1}x_{1}+O(1).$
Lemma 36.
For any $i>1$, we have $x_{i}=\frac{\lambda+w_{1}+2e(W_{1})/\lambda}{\lambda+w_{i}}x_{1}+O(n^{-2})$.
Proof.
We have
$\lambda x_{i}$ $=\lambda x_{1}-w_{i}x_{i}+X_{1}$
$=\lambda x_{1}-w_{i}x_{i}+w_{1}x_{1}+\frac{2e(W_{1})}{\lambda}x_{1}+O(n^{-1})$
and solving for $x_{i}$ gives the result. ∎
Lemma 37.
We have $w_{i}=n/r+O(1)$ for each $i$.
Proof.
It suffices to show that for each $i>1$, we have $|w_{1}-w_{i}|=O(1)$. Taking Equation 1 with respect to $G_{1i}$ gives
00 0 $\geq w_{1}x_{1}-w_{i}x_{i}+O(1)$
$=w_{1}x_{1}-w_{i}\frac{\lambda+w_{1}+O(1)}{\lambda+w_{i}}x_{1}+O(1)=\left(w_{1}-w_{i}\frac{\lambda+w_{i}}{\lambda+w_{1}}\right)x_{1}$
$O(n)$ $\geq(\lambda+w_{i})w_{1}-(\lambda+w_{1})w_{i}=\lambda(w_{1}-w_{i})$
from which we obtain $w_{1}-w_{i}\leq O(1)$. Similarly, taking Equation 1 with respect to $G_{i1}$ gives
00 0 $\geq w_{i}x_{i}-w_{1}x_{1}+O(1)$
$=w_{i}\frac{\lambda+w_{1}+O(1)}{\lambda+w_{i}}x_{1}-w_{1}x_{1}+O(1)$
$=\left(w_{i}\frac{\lambda+w_{1}}{\lambda+w_{i}}-w_{1}\right)x_{1}+O(1)$
$O(n)$ $\geq(\lambda+w_{1})w_{i}-(\lambda+w_{i})w_{1}=\lambda(w_{i}-w_{1})$
from which we obtain $w_{i}-w_{1}\leq O(1)$. ∎
Combining 36 and 37 now gives $x_{i}=x_{1}+O(n^{-1})$ for every $i$. Furthermore, combining 37 with the bound $\lambda\leq\Delta(G)$ proves $\lambda\leq\frac{r-1}{r}n+O(1)$, hence $\lambda=\frac{r-1}{r}n+O(1)$. These estimates will be useful in the rest of the argument.
Lemma 38.
We have $|w_{i}-w_{j}|\leq 1$ for each $i,j>1$.
Proof.
If there is some $i,j>1$ such that $w_{j}\geq w_{i}+2$, then taking Equation 1 with respect to $G_{ji}$ gives
$0\geq$ $\sum_{v\in W_{j}-u}x_{v}-\sum_{v\in W_{i}}x_{v}$
$=$ $(w_{j}-1-|W_{j}^{m}|)\left(\frac{\lambda+w_{i}}{\lambda+w_{j}}x_{i}+O(n^{-2})\right)+|W_{j}^{m}|\left(\frac{\lambda+w_{i}}{\lambda+w_{j}}x_{i}+O(n^{-1})\right)$
$-(w_{i}-|W_{i}^{m}|)x_{i}-|W_{i}^{m}|(x_{i}+O(n^{-1}))$
$=$ $\left((w_{j}-1)\frac{\lambda+w_{i}}{\lambda+w_{j}}-w_{i}\right)x_{i}+O(n^{-1}).$
Now since $w_{i}<w_{j}$, we have $w_{i}<n/2$, and since $r\geq 3$ we have $\lambda\geq\left(1-\frac{1}{r}\right)n+O(1)\geq 2n/3+O(1)$. This implies that $\lambda>w_{i}+n^{1/2}$, which together with $w_{j}\geq w_{i}+2$ implies $\lambda(w_{j}-w_{i}-1)>w_{i}+n^{1/2}$, so $(w_{j}-1)\frac{\lambda+w_{i}}{\lambda+w_{j}}-w_{i}\geq\Omega(n^{-1/2})$ and so the expression displayed above is positive, a contradiction. ∎
Lemma 39.
For all $i>1$, we have $w_{1}\leq w_{i}+2$.
Proof.
Otherwise, we have $w_{1}\geq w_{i}+3$. Then taking Equation 1 with respect to $G_{1i}$, where $u$ is pendant in $G[W_{1}]$, gives
00 0 $\geq w_{1}x_{1}+\frac{2e(W_{1})}{\lambda}x_{1}-2x_{1}-w_{i}x_{i}+O(n^{-1})$
$=w_{1}x_{1}-2+\frac{2e(W_{1})}{\lambda}x_{1}-w_{i}\frac{\lambda+w_{1}+2e(W_{1})/\lambda}{\lambda+w_{i}}x_{1}+O(n^{-1})$
$=(w_{1}-w_{i})\left(1-\frac{w_{i}}{\lambda+w_{i}}\right)x_{1}-2+2e(W_{1})\left(\frac{1}{\lambda}-\frac{w_{i}}{\lambda(\lambda+w_{i})}\right)x_{1}+O(n^{-1})$
$\geq 3\left(1-\frac{w_{i}}{\lambda+w_{i}}\right)-2+\frac{2e(W_{1})}{\lambda}\left(1-\frac{w_{i}}{\lambda+w_{i}}\right)+O(n^{-1})$
$=3\left(1-\frac{1}{r}\right)-2+\frac{2e(W_{1})}{\lambda}\left(1-\frac{w_{i}}{\lambda+w_{i}}\right)+O(n^{-1})>0.$
Now if $e(W_{1})=o(n)$ then we must have $k(\mathcal{F})=1$ and Theorem 1 gives $G\in\mathrm{EX}(n,\mathcal{F})$. Otherwise, we have $\frac{2e(W_{1})}{\lambda}\geq\frac{1}{r-1}+o(1)$ and $1-\frac{w_{i}}{\lambda+w_{i}}=1-\frac{1}{r}+o(1)$, so that
$00\geq 3\left(1-\frac{1}{r}\right)-2+\frac{1}{r}+o(1)\geq 1-\frac{2}{r}+o(1)>$ 0
a contradiction. ∎
Lemma 40.
We have $w_{i}\leq w_{1}+2$ for all $i$.
Proof.
Otherwise we have $w_{i}\geq w_{1}+3$. Then taking Equation 1 with respect to $G_{i1}$ gives
00 0 $\geq(w_{i}-1)x_{i}-w_{1}x_{1}-\frac{2e(W_{1})}{\lambda}x_{1}+O(n^{-1})$
$=w_{i}\frac{\lambda+w_{1}+2e(W_{1})/\lambda}{\lambda+w_{i}}x_{1}-1-w_{1}x_{1}-\frac{2e(W_{1})}{\lambda}+O(n^{-1})$
$=w_{i}\left(1+\frac{w_{1}-w_{i}}{\lambda+w_{i}}\right)x_{1}-w_{1}x_{1}-1+2e(W_{1})\left(\frac{w_{i}}{\lambda(\lambda+w_{i})}-\frac{1}{\lambda}\right)+O(n^{-1})$
$=(w_{i}-w_{1})\left(1-\frac{w_{i}}{\lambda+w_{i}}\right)-1+\frac{2e(W_{1})}{\lambda}\left(\frac{w_{i}}{\lambda+w_{i}}-1\right)+O(n^{-1}).$
Now using $e(W_{1})\leq\frac{n}{r}+O(1)$ we obtain $\frac{2e(W_{1})}{\lambda}\left(\frac{w_{i}}{\lambda+w_{i}}-1\right)\geq-\frac{2}{r}+o(1)$, so
$00\geq 3\left(1-\frac{1}{r}\right)-1-\frac{2}{r}+o(1)\geq 2-\frac{5}{3}+o(1)>$ 0
a contradiction. ∎
We now state a general lemma which will be useful for the rest of the proof.
Lemma 41.
Suppose $H$ is any graph obtained from $K_{W_{1},\ldots,W_{r}}$ by embedding a graph of bounded degree in $W_{1}$, then deleting and adding $O(1)$ edges. Assume that $|W_{i}|=n/r+O(1)$ for each $i$. If $H^{\prime}$ is obtained from $H$ by deleting $a$ edges and adding $b$ edges where $a,b=O(1)$, then $\lambda(H^{\prime})-\lambda(H)=2(b-a)/n+O(n^{-2})$; in particular if $a<b$ then $\lambda(H^{\prime})>\lambda(H)$.
Proof.
Let $x$ be the principal eigenvector of $H$, scaled so that its largest entry is $1$. From the estimates proved above, we already have that $x_{v}=1+O(n^{-1})$ for every $v$. Let $A^{\prime}$ be the set of deleted edges and $B^{\prime}$ be the set of added edges. Then we have
$x^{T}A(H^{\prime})x-x^{T}A(H)x$ $=2\sum_{uv\in B^{\prime}}x_{u}x_{v}-2\sum_{uv\in A^{\prime}}x_{u}x_{v}$
$=2b(1+O(n^{-1}))^{2}-2a(1+O(n^{-1}))^{2}=2(b-a)+O(n^{-1}).$
Thus
$\lambda(H^{\prime})-\lambda(H)$ $\geq\frac{x^{T}A(H^{\prime})x}{x^{T}x}-\frac{x^{T}A(H)x}{x^{T}x}$
$=\frac{2(b-a)+O(n^{-1})}{n+O(1)}=\frac{2(b-a)}{n}+O(n^{-2}).$
By symmetry, we also have
$\lambda(H)-\lambda(H^{\prime})\geq\frac{2(a-b)}{n}+O(n^{-2})$
which proves the result. ∎
Lemma 42.
For each $i=2,\ldots,r$, we have $w_{i}\leq w_{1}+1$.
Proof.
If $w_{i}=w_{1}+2$, then as in 40, taking Equation 1 with respect to $G_{i1}$ gives
$0\geq 2\left(1-\frac{1}{r}\right)-1+\frac{2e(W_{1})}{\lambda}\left(\frac{w_{i}}{\lambda+w_{i}}-1\right)+O(n^{-1}).$
We have $2e(W_{1})\leq\frac{2(k-1)}{k}\frac{n}{r}+O(1)$, so we obtain
$0\geq 1-\frac{2}{r}-\frac{k-1}{k}\frac{2}{r}+o(1)$
which gives a contradiction if $r\geq 4$. If $r=3$, we require a more detailed argument. Without loss of generality, $i=2$. Let $G^{\prime}$ be obtained from $G$ by converting every vertex in $W^{m}$ to an $W$-ordinary vertex in the same part, let $a=e(G^{\prime})-e(G)=O(1)$, and let $G_{21}^{\prime}=(G^{\prime})_{21}$. Observe by 41 that $\lambda(G^{\prime})=\lambda(G)+2a/n+O(n^{-2})$ and $\lambda(G_{21})=\lambda(G_{21}^{\prime})-2a/n+O(n^{-2})$, and so it suffices to show that $\lambda(G_{21}^{\prime})-\lambda(G^{\prime})\geq\delta/n$ for some constant $\delta>0$. Let $V(G_{21}^{\prime})$ have partition $U_{1}\sqcup U_{2}\sqcup U_{3}$ so that $|U_{1}|=w_{1}+1$, $|U_{2}|=w_{2}-1$, and $|U_{3}|=w_{3}$. Note that according to the previous lemmas, we either have $w_{1}=w_{2}-2=w_{3}-1$ or $w_{1}=w_{2}-2=w_{3}-2$. Let $\mu=\lambda(G^{\prime})$, $\rho=\lambda(G_{21}^{\prime})$, with corresponding eigenvectors $y,z$, respectively, scaled so that $y_{1}:=\frac{1}{\mu}\sum_{v\not\in W_{1}}x_{v}=1$ and $z_{1}:=\frac{1}{\rho}\sum_{v\not\in U_{1}}z_{v}=1$. Let $y_{2},y_{3}$ be the common eigenweights of vertices in $W_{2},W_{3}$, and let $z_{2},z_{3}$ be the common eigenweights of vertices in $U_{2},U_{3}$.
Claim 14.
We have $\rho-\mu=O(n^{-1})$,
Proof.
Observe that $G_{21}^{\prime}$ can be obtained from $G^{\prime}$ by deleting $w_{1}$ edges and adding $w_{2}$ edges, so
$\frac{1}{2}(\lambda(G_{21}^{\prime})-\lambda(G^{\prime}))\geq\frac{-w_{1}(1+O(n^{-1}))^{2}+w_{2}(1+O(n^{-1}))^{2}}{y^{T}y}=\frac{O(1)}{y^{T}y}=O(n^{-1}).$
The reverse direction is proved similarly. ∎
Now for $v\in W_{1}$ we obtain
$\mu^{2}y_{v}$ $=\mu\left(\mu y_{1}+\sum_{u\in N_{W_{1}}(v)}x_{u}\right)$
$=\mu^{2}+\mu d_{W_{1}}(v)+d_{W_{1}}^{\prime}(v)+O(n^{-1})$
$y_{v}$ $=1+\frac{d_{W_{1}}(v)}{\mu}+\frac{d_{W_{1}}^{\prime}(v)}{\mu^{2}}+O(n^{-3}).$
Similarly for $v\in W_{1}$,
$z_{v}$ $=1+\frac{d_{W_{1}}(v)}{\rho}+\frac{d_{W_{1}}^{\prime}(v)}{\rho^{2}}+O(n^{-3})$
$=1+\frac{d_{W_{1}}(v)}{\mu}+\frac{d_{W_{1}}^{\prime}(v)}{\mu^{2}}+O(n^{-3}).$
and for the vertex $v\in U_{1}-W_{1}$ we have $z_{v}=1$. Let $Y_{1}=\sum_{v\in W_{1}}y_{v}$, $Z_{1}=\sum_{v\in U_{1}}z_{v}$. Then using the eigenweight estimates above we obtain
$Y_{1}$ $=w_{1}+\frac{2e(W_{1})}{\mu}+\frac{\sum_{v\in W_{1}}d_{W_{1}}^{\prime}(v)}{\mu^{2}}+O(n^{-2})$
$Z_{1}$ $=w_{1}+1+\frac{2e(W_{1})}{\mu}+\frac{\sum_{v\in W_{1}}d_{W_{1}}^{\prime}(v)}{\mu^{2}}+O(n^{-2})=Y_{1}+1+O(n^{-2}).$
In particular, $Y_{1},Z_{1}=w_{1}+O(1)$. From the first-degree eigenvector equations we obtain
$y_{2}$ $=1+\frac{Y_{1}-w_{2}}{\mu+w_{2}}$
$y_{3}$ $=1+\frac{Y_{1}-w_{3}}{\mu+w_{3}}$
$z_{2}$ $=1+\frac{Z_{1}-w_{2}+1}{\rho+w_{2}-1}=1+\frac{Y_{2}-w_{2}+2}{\mu+w_{2}-1}+O(n^{-3})$
$=1+\frac{Y_{2}-w_{2}+2}{\mu+w_{2}}+\frac{Y_{2}-w_{2}+2}{(\mu+w_{2})(\mu+w_{2}-1)}+O(n^{-3})$
$z_{3}$ $=1+\frac{Z_{1}-w_{3}}{\rho+w_{3}}=1+\frac{Y_{1}-w_{3}+1}{\mu+w_{3}}+O(n^{-3}).$
Then
$\mu=$ $\,\mu y_{1}=w_{2}\left(1+\frac{Y_{1}-w_{2}}{\mu+w_{2}}\right)+w_{3}\left(1+\frac{Y_{1}-w_{3}}{\mu+w_{3}}\right)$ (2)
$\rho=$ $\,\rho z_{1}=(w_{2}-1)\left(1+\frac{Y_{1}-w_{2}+2}{\mu+w_{2}}+\frac{Y_{1}-w_{2}+2}{(\mu+w_{2})(\mu+w_{2}-1)}\right)$
$+w_{3}\left(1+\frac{Y_{1}-w_{3}+1}{\mu+w_{3}}\right)+O(n^{-2})$
hence
$\rho-\mu=$ $-1+\frac{3w_{2}-Y_{1}-2}{\mu+w_{2}}+\frac{(w_{2}-1)(Y_{1}-w_{2}+2)}{(\mu+w_{2})(\mu+w_{2}-1)}+\frac{w_{3}}{\mu+w_{3}}+O(n^{-2})$
Now $Y_{1}-w_{2}=\frac{2e(W_{1})}{\mu}-2+O(n^{-1})=\frac{k-1}{k}-2+O(n^{-1})$, so
$\rho-\mu$ $=-1+\frac{2w_{2}-\frac{k-1}{k}}{\mu+w_{2}}+\frac{w_{2}\frac{k-1}{k}}{(\mu+w_{2})(\mu+w_{2}-1)}+\frac{w_{3}}{\mu+w_{3}}+O(n^{-2})$
$=-1+\frac{2w_{2}}{\mu+w_{2}}-\frac{2(k-1)}{3k}\frac{1}{n}+\frac{w_{3}}{\mu+w_{3}}+O(n^{-2}).$
Case 1: $w_{3}=w_{2}-1$. Using Equation 2 we obtain an initial estimate of $\mu$:
$\mu$ $=w_{2}+w_{3}+\frac{w_{2}(Y_{1}-w_{2})}{\mu+w_{2}}+\frac{(w_{2}-1)(Y_{1}-w_{2}+1)}{\mu+w_{3}}$
$=2w_{2}-1+\frac{w_{2}(2e(W_{1})/\mu-2)}{\mu+w_{2}}+\frac{w_{2}(2e(W_{1})/\mu-1)}{\mu+w_{2}-1}+O(n^{-1})$
$=2w_{2}-1+\frac{w_{2}(4e(W_{1})/\mu-3)}{\mu+w_{2}}+O(n^{-1})$
$=2w_{2}-2+\frac{2}{3}\frac{k-1}{k}+O(n^{-1}).$
Also, note that
$\frac{w_{3}}{\mu+w_{3}}$ $=\frac{w_{2}-1}{\mu+w_{2}-1}=\frac{w_{2}}{\mu+w_{2}}+\frac{w_{2}}{(\mu+w_{2})(\mu+w_{2}-1)}-\frac{1}{n}+O(n^{-2})$
$=\frac{w_{2}}{\mu+w_{2}}-\frac{2}{3n}+O(n^{-2}).$
Thus
$\rho-\mu$ $=-1+\frac{\mu}{\mu+w_{2}}+\frac{w_{2}}{\mu+w_{2}}+\left(2-\frac{2}{3}\frac{k-1}{k}-\frac{2}{3}\frac{k-1}{k}-\frac{2}{3}\right)\frac{1}{n}+O(n^{-2})$
$=\frac{4}{3}\left(1-\frac{k-1}{k}\right)\frac{1}{n}+O(n^{-2})$
and we are done.
Case 2: $w_{3}=w_{2}$. Similarly to case 1 we obtain an initial estimate of $\mu$:
$\mu$ $=w_{2}+w_{3}+\frac{w_{2}(Y_{1}-w_{2})}{\mu+w_{2}}+\frac{w_{2}(Y_{1}-w_{2})}{\mu+w_{2}}$
$=2w_{2}+\frac{2w_{2}(Y_{1}-w_{2})}{\mu+w_{2}}$
$=2w_{2}-\frac{4}{3}+\frac{2}{3}\frac{k-1}{k}+O(n^{-1}).$
Thus,
$\rho-\mu$ $=-1+\frac{\mu}{\mu+w_{2}}+\frac{w_{2}}{\mu+w_{2}}+\left(\frac{4}{3}-\frac{2}{3}\frac{k-1}{k}-\frac{2}{3}\frac{k-1}{k}\right)\frac{1}{n}+O(n^{-2})$
$=\frac{4}{3}\left(1-\frac{k-1}{k}\right)\frac{1}{n}+O(n^{-2})$
and we are done. ∎
Henceforth, whenever we need precise eigenvalue estimates we will follow the argument of 42, so we will not repeat every detail of such arguments. Before completing our analysis of the part sizes of $G$, we need the following lemma.
Lemma 43.
Suppose $G_{1}$, $G_{2}$ are graphs obtained from $T_{n,r}$ by embedding the same number of trees of order $k\geq 4$ in a large part so as to cover all but $O(1)$ vertices in the part. Then
$\lambda(G_{1})-\lambda(G_{2})\leq\left[\frac{k-5+6/k}{r-1}\right]\frac{1}{n}+O(n^{-2}).$
Moreover, if all the trees embedded in $G_{1}$ are stars and all the trees embedded in $G_{2}$ are paths, then equality holds.
Proof.
Let $G_{\mathrm{star}}$ and $G_{\mathrm{path}}$ be the graphs obtained from $G_{1}$ or $G_{2}$ by replacing each embedded tree with a star or path, respectively.
Claim 15.
Let $K$ be any graph obtained by embedding a disjoint union of trees of order $k$ in one part of $T_{n,r}$, where one of the trees $T$ is not a star. Then there is another tree $T^{\prime}$ with more pendant vertices than $T$ such that, if $K^{\prime}$ is obtained from $K$ by replacing $T$ with $T^{\prime}$, we have $\lambda(K^{\prime})\geq\lambda(K)$.
Proof.
Choose a non-pendant vertex $u$ as the root of $T$ and choose $v\in V(T)$ at maximum distance from $u$. Since $T$ is not a star, $d(v,u)\geq 2$ and also $v$ is a pendant vertex. Let $v\sim_{T}w$. Since $v$ is at maximum distance from $u$ and $w\neq u$, $w$ has a unique neighbor $w^{\prime}$ such that $w^{\prime}$ is not a pendant vertex. Let $T^{\prime}$ be obtained by deleting the edge $wv^{\prime}$ and adding the edge $w^{\prime}v^{\prime}$ for every pendant $v^{\prime}\sim_{T}w^{\prime}$, and note that $T^{\prime}$ has more pendant vertices than $T$. Now $K^{\prime}$ is obtained from $K$ by performing a Kelmans transformation and so $\lambda(K^{\prime})\geq\lambda(K)$ (see e.g. [csikvari2009conjecture]). ∎
By repeatedly applying 15 to $G_{1}$, it follows that $\lambda(G_{1})\leq\lambda(G_{\mathrm{star}})$.
Claim 16.
Let $K$ be any graph obtained by embedding a disjoint union of trees of order $k$ in one part of $T_{n,r}$, where one of the trees $T$ is not a path. Then there is another tree $T^{\prime}$ of greater diameter than $T$ such that, if $K^{\prime}$ is obtained from $K$ by replacing $T$ with $T^{\prime}$, we have $\lambda(K^{\prime})\leq\lambda(K)$.
Proof.
Let a longest path $P$ in $T$ have endpoints $u$ and $u^{\prime}$ (so that $u,u^{\prime}$ are pendant in $T$), and note that if $T$ is not a path then there is a vertex $v$ of degree at least $3$ which is on $P$. Let $v\sim w$ with $w\not\in\{u,u^{\prime}\}$. Let $T^{\prime}$ be obtained from $T$ by deleting $vw$ and adding $uw$ and note that $T^{\prime}$ is a tree of greater diameter than $T$. Let $\mu=\lambda(K)$, with eigenvector $y$, scaled so that the sum of $y_{v}$ over $v$ not in the large part equals $\mu$. For any $x\in V(T)$ one can show $y_{x}=1+\frac{d_{T}(x)}{\mu}+O(n^{-2})$. Therefore
$\frac{1}{2y_{w}}(y^{T}A(K^{\prime})y-y^{T}A(K)y)$ $\leq 1+\frac{1}{\mu}-1-\frac{3}{\mu}+O(n^{-2})<0.$
By repeatedly applying 16 to $G_{2}$, it follows that $\lambda(G_{2})\geq\lambda(G_{\mathrm{path}})$. Let $V_{1},\ldots,V_{r}$ be the common partition of $G_{\mathrm{star}}$ and $G_{\mathrm{path}}$, with $v_{i}:=|V_{i}|$. Let $v_{1},\ldots,v_{h}=v_{1}$ and $v_{h+1},\ldots,v_{r}=v_{1}-1$. Let $\mu=\lambda(G_{\mathrm{star}})$, $\rho=\lambda(G_{\mathrm{path}})$, and let $y,z$ be the corresponding eigenvectors, scaled so that $y_{1}:=\frac{1}{\mu}\sum_{v\not\in V_{1}}y_{v}=1$ and $z_{1}:=\frac{1}{\rho}\sum_{v\not\in V_{1}}z_{v}=1$. Similarly to 14, we find $\mu-\rho=O(n^{-1})$. For $v\in V_{1}$, let $d_{s}(v):=d_{G_{\mathrm{star}}[V_{1}]}(v)$ and $d_{p}(v):=d_{G_{\mathrm{path}[V_{1}]}}(v)$, and let $d_{s}^{\prime}(v)$ ($d_{p}^{\prime}(v)$) be the number of walks of length 2 starting from $v$ in $G_{\mathrm{star}}[V_{1}]$ ($G_{\mathrm{path}}[V_{1}]$). The second-degree eigenvector equations give that for $v\in V_{1}$,
$y_{v}$ $=1+\frac{d_{s}(v)}{\mu}+\frac{d_{s}^{\prime}(v)}{\mu^{2}}+O(n^{-3})$
$z_{v}$ $=1+\frac{d_{p}(v)}{\mu}+\frac{d_{p}^{\prime}(v)}{\mu^{2}}+O(n^{-3}).$
Let $Y_{1}=\sum_{v\in V_{1}}y_{v}$, $Z_{1}=\sum_{v\in V_{1}}z_{v}$. We obtain
$Y_{1}$ $=v_{1}+\frac{2e(V_{1})}{\mu}+\frac{1}{\mu^{2}}\sum_{v\in V_{1}}d_{s}^{\prime}(v)+O(n^{-2})$
$Z_{1}$ $=v_{1}+\frac{2e(V_{1})}{\mu}+\frac{1}{\mu^{2}}\sum_{v\in V_{1}}d_{p}^{\prime}(v)+O(n^{-2}).$
Observe that the number of walks of length two in a path of order $k\geq 4$ is $64k-$ 6 and the number of walks of length two in $K_{1,k-1}$ is $k(k-1)$. Therefore,
$Y_{1}-Z_{1}$ $=\frac{1}{\mu^{2}}\sum_{v\in V_{1}}(d_{s}^{\prime}(v)-d_{p}^{\prime}(v))+O(n^{-2})$
$=\frac{1}{\mu^{2}}\left(\frac{v_{1}-O(1)}{k}k(k-1)-\frac{v_{1}-O(1)}{k}(4k-6)\right)+O(n^{-2})$
$=\frac{v_{1}(k-5+6/k)}{\mu^{2}}+O(n^{-2}).$
Now for $v\in V_{i}$ $(i\geq 2)$, the eigenvector equations give
$y_{v}=\begin{cases}1+\frac{Y_{1}-v_{1}}{\mu+v_{1}},&2\leq i\leq h\\ 1+\frac{Y_{1}-v_{1}+1}{\mu+v_{1}}+\frac{Y_{1}-v_{1}+1}{(\mu+v_{1})^{2}}+O(n^{-3}),&h+1\leq i\leq r\end{cases}$
and
$z_{v}=\begin{cases}1+\frac{Z_{1}-v_{1}}{\mu+v_{1}}+O(n^{-3}),&2\leq i\leq h\\ 1+\frac{Z_{1}-v_{1}+1}{\mu+v_{1}}+\frac{Z_{1}-v_{1}+1}{(\mu+v_{1})^{2}}+O(n^{-3}),&h+1\leq i\leq r.\end{cases}$
We obtain
$\mu=$ $\,\mu y_{1}=(h-1)v_{1}\left(1+\frac{Y_{1}-v_{1}}{\mu+v_{1}}\right)$
$+(r-h)(v_{1}-1)\left(1+\frac{Y_{1}-v_{1}+1}{\mu+v_{1}}+\frac{Y_{1}-v_{1}+1}{(\mu+v_{1})^{2}}\right)+O(n^{-2})$
$\rho=$ $\,\rho z_{1}=(h-1)v_{1}\left(1+\frac{Z_{1}-v_{1}}{\mu+v_{1}}\right)$
$+(r-h)(v_{1}-1)\left(1+\frac{Z_{1}-v_{1}+1}{\mu+v_{1}}+\frac{Z_{1}-v_{1}+1}{(\mu+v_{1})^{2}}\right)+O(n^{-2}).$
Thus
$\mu-\rho$ $=(h-1)v_{1}\frac{Y_{1}-Z_{1}}{\mu+v_{1}}+(r-h)(v_{1}-1)\frac{Y_{1}-Z_{1}}{\mu+v_{1}}+(r-h)(v_{1}-1)\frac{Y_{1}-Z_{1}}{(\mu+v_{1})^{2}}+O(n^{-2})$
$=\frac{(r-1)v_{1}(Y_{1}-Z_{1})}{\mu+v_{1}}+O(n^{-2})$
$=\left[\frac{k-5+6/k}{r-1}\right]\frac{1}{n}+O(n^{-2}).$
This proves the equality part of the result, and the inequality part follows from $\lambda(G_{\mathrm{path}})\leq\lambda(G_{1}),\lambda(G_{2})\leq\lambda(G_{\mathrm{star}}).$
Lemma 44.
For each $i=2,\ldots,r$, we have $w_{i}\geq w_{1}-1$.
Proof.
Suppose not, and assume without loss of generality that $w_{1}=w_{2}+2$. Let $w_{2},\ldots,w_{h}=w_{1}-2$ and $w_{h+1},\ldots,w_{r}=w_{1}-1$. Let $G^{\prime}$ be obtained from $G$ by converting every vertex in $W^{m}$ to a $W$-ordinary vertex, and let $a=e(G^{\prime})-e(G)=O(1)$. Let $G^{\prime\prime}$ be obtained from $G^{\prime}$ by deleting a pendant edge in $G^{\prime}[W_{1}]$. Then $e(G^{\prime\prime})=e(G)+a-1$, and $G^{\prime\prime}$ has an isolated vertex $u^{\prime}$ in $G^{\prime\prime}[W_{1}]$. Let $\mu=\lambda(G^{\prime\prime})$, $\rho=\lambda(G_{12}^{\prime\prime})$, where $G_{12}^{\prime\prime}$ is obtained by transferring $u^{\prime}$ to $W_{2}$; then $G_{12}^{\prime\prime}$ is $\mathcal{F}$-free and $e(G_{12}^{\prime\prime})=e(G^{\prime\prime})+1=e(G)+a$. Let $y$ be the principal eigenvector of $G^{\prime\prime}$ and $z$ be the principal eigenvector of $G_{12}^{\prime\prime}$, scaled so that $y_{1}:=\frac{1}{\mu}\sum_{v\not\in W_{1}}y_{v}=1$ and $z_{1}:=\frac{1}{\rho}\sum_{v\not\in W_{1}-\{u^{\prime}\}}z_{v}=1$. Similarly to 14, we find $\rho-\mu=O(n^{-1})$.
Claim 17.
$\rho-\mu=\left[2-\frac{2}{r}+\frac{4(k-1)}{kr}\right]\frac{1}{n}+O(n^{-2}).$
Proof.
For $v\in W_{1}$ the second-degree eigenvector equation gives
$y_{v}=1+\frac{d_{W_{1}}(v)}{\mu}+\frac{d_{W_{1}}^{\prime}(v)}{\mu^{2}}+O(n^{-3})$
and for $v\in W_{1}-\{u^{\prime}\}$, gives
$z_{v}=1+\frac{d_{W_{1}}(v)}{\mu}+\frac{d_{W_{1}}^{\prime}(v)}{\mu^{2}}+O(n^{-3}).$
Let $Y_{1}=\sum_{v\in W_{1}}y_{v}$, $Z_{1}=\sum_{v\in W_{1}-\{u^{\prime}\}}z_{v}$. Then we obtain
$Y_{1}$ $=w_{1}+\frac{2(e_{G}(W_{1})-1)}{\mu}+\frac{\sum_{v\in W_{1}-\{u^{\prime}\}}d_{W_{1}}^{\prime}(v)}{\mu^{2}}+O(n^{-2})$
$Z_{1}$ $=Y_{1}-1+O(n^{-2}).$
In particular $Y_{1}=w_{1}+\frac{2(k-1)}{k(r-1)}+O(n^{-1})$. The first-degree eigenvector equations now give that if $v\in W_{i}$ with $i\geq 2$, we have
$y_{v}=\begin{cases}1+\frac{Y_{1}-w_{2}}{\mu+w_{2}},&2\leq i\leq h\\ 1+\frac{Y_{1}-w_{2}-1}{\mu+w_{2}}-\frac{Y_{1}-w_{2}-1}{(\mu+w_{2})^{2}}+O(n^{-3}),&i>h\end{cases}$
and if $v\in W_{2}\cup\{u^{\prime}\}$ or $v\in W_{i}$ for $i>2$, then
$z_{v}=\begin{cases}1+\frac{Y_{1}-w_{2}-1}{\mu+w_{2}}+O(n^{-3}),&3\leq i\leq h\\ 1+\frac{Y_{1}-w_{2}-2}{\mu+w_{2}}-\frac{Y_{1}-w_{2}-2}{(\mu+w_{2})^{2}}+O(n^{-3}),&i=2\text{ or }i>h.\end{cases}$
Thus, $\mu=\mu y_{1}$ and $\rho=\rho z_{1}$ give
$\mu=$ $\,(h-1)w_{2}\left(1+\frac{Y_{1}-w_{2}}{\mu+w_{2}}\right)$
$+(r-h)(w_{2}+1)\left(1+\frac{Y_{1}-w_{2}-1}{\mu+w_{2}}-\frac{Y_{1}-w_{2}-1}{(\mu+w_{2})^{2}}\right)+O(n^{-2})$
$\rho=$ $\,(h-2)w_{2}\left(1+\frac{Y_{1}-w_{2}-1}{\mu+w_{2}}\right)$
$+(r-h+1)(w_{2}+1)\left(1+\frac{Y_{1}-w_{2}-2}{\mu+w_{2}}-\frac{Y_{1}-w_{2}-2}{(\mu+w_{2})^{2}}\right)+O(n^{-2}).$
We now obtain an initial estimate on $\mu$. We have
$\mu=$ $\,(h-1)w_{2}+(r-h)(w_{2}+1)+\frac{(h-1)w_{2}\left(w_{1}-w_{2}+\frac{2(k-1)}{k(r-1)}\right)+O(1)}{\mu+w_{2}}$
$+\frac{(r-h)(w_{2}+1)\left(w_{1}-w_{2}-1+\frac{2(k-1)}{k(r-1)}\right)+O(1)}{\mu+w_{2}}+O(n^{-1})$
$=$ $\,(r-1)w_{2}+r-h+\frac{h}{r}-\frac{2}{r}+1+\frac{2(k-1)}{kr}+O(n^{-1}).$
Now
$\rho-\mu=$ $\,-w_{2}+w_{2}\frac{(h-2)(Y_{1}-w_{2}-1)-(h-1)(Y_{1}-w_{2})}{\mu+w_{2}}$
$+w_{2}+1+(w_{2}+1)\frac{(r-h+1)(Y_{1}-w_{2}-2)-(r-h)(Y_{1}-w_{2}-1)}{\mu+w_{2}}$
$-(w_{2}+1)\frac{(r-h+1)(Y_{1}-w_{2}-2)-(r-h)(Y_{1}-w_{2}-1)}{(\mu+w_{2})^{2}}+O(n^{-2})$
$=$  $\,1+\frac{w_{2}(w_{2}-Y_{1}-h+2)}{\mu+w_{2}}$
$+\frac{(w_{2}+1)(Y_{1}-w_{2}-r+h-2)}{\mu+w_{2}}$
$-\frac{(w_{2}+1)(Y_{1}-w_{2}-r+h-2)}{(\mu+w_{2})^{2}}+O(n^{-2})$
$=$  $\,1-\frac{rw_{2}}{\mu+w_{2}}+\frac{Y_{1}-w_{2}-r+h-2}{\mu+w_{2}}-\frac{w_{2}(Y_{1}-w_{2}-r+h-2)}{(\mu+w_{2})^{2}}+O(n^{-2}).$
We estimate each term. Using the initial estimates of $\mu$ and $Y_{1}$, we have
$\frac{rw_{2}}{\mu+w_{2}}$ $=1-\left[r-h+\frac{h}{r}-\frac{2}{r}+1+\frac{2(k-1)}{kr}\right]\frac{1}{n}+O(n^{-2})$
$\frac{Y_{1}-w_{2}-r+h-2}{\mu+w_{2}}$ $=\left[\frac{2(k-1)}{k(r-1)}-r+h\right]\frac{1}{n}+O(n^{-2})$
$\frac{w_{2}(Y_{1}-w_{2}-r+h-2)}{(\mu+w_{2})^{2}}$ $=\left[\frac{2(k-1)}{kr(r-1)}-1+\frac{h}{r}\right]\frac{1}{n}+O(n^{-1}).$
So,
$\rho-\mu$ $=\left[-\frac{2}{r}+1+\frac{2(k-1)}{kr}+\frac{2(k-1)}{k(r-1)}-\frac{2(k-1)}{kr(r-1)}+1\right]\frac{1}{n}+O(n^{-2})$
$=\left[2-\frac{2}{r}+\frac{4(k-1)}{kr}\right]\frac{1}{n}+O(n^{-2}).$
Note that $G_{12}^{\prime\prime}$ is obtained from $T_{n,r}$ by embedding trees of order $k$ in a large part, with possible 1 smaller tree. By 1 and applying a vertex transfer if necessary, some graph $H\in\mathrm{EX}(n,\mathcal{F})$ is obtained from $T_{n,r}$ by embedding trees of order $k$ in a large part, then making a bounded number of edge modifications. There is a graph $H^{\prime}$ obtained from $H$ by a bounded number of edge modifications and a graph $G_{12}^{\prime\prime\prime}$ obtained from $G_{12}^{\prime\prime}$ by a bounded number of edge modifications, so that each graph is obtained from $T_{n,r}$ by embedding the same number of trees of order $k$ in the large part; let $b=e(H)-e(H^{\prime})=O(1)$, $a^{\prime}=e(G_{12}^{\prime\prime\prime})-e(G_{12}^{\prime\prime})=O(1).$ Now if $G\in\mathrm{EX}(n,\mathcal{F})$ we are done, so we may assume $b+a^{\prime}+a=e(H)-e(G)\geq 1$.
Case 1: $k\in\{1,2,3\}$. Then there is only one tree of order $k$, so $H^{\prime}$ can be obtained from $G_{12}^{\prime\prime\prime}$ by a bounded number of edge modifications. By 41, we have
$\lambda(H)-\lambda(G)=$ $\,\lambda(H)-\lambda(H^{\prime})+\lambda(H^{\prime})-\lambda(G_{12}^{\prime\prime\prime})+\lambda(G_{12}^{\prime\prime\prime})-\lambda(G_{12}^{\prime\prime})$
$+\lambda(G_{12}^{\prime\prime})-\lambda(G^{\prime\prime})+\lambda(G^{\prime\prime})-\lambda(G)$
$=$ $\,\left[2b+0+2a^{\prime}+2-\frac{2}{r}+\frac{4(k-1)}{kr}+2(a-1)\right]\frac{1}{n}+O(n^{-2})>0$
which gives a contradiction.
Case 2: $k\geq 4$. By 43, we have $\lambda(H^{\prime})-\lambda(G_{12}^{\prime\prime\prime})\geq-\left[\frac{k-5+6/k}{r-1}\right]\frac{1}{n}+O(n^{-2})$. Then by 41 we have
$\lambda(H)-\lambda(G)=$ $\,\lambda(H)-\lambda(H^{\prime})+\lambda(H^{\prime})-\lambda(G_{12}^{\prime\prime\prime})+\lambda(G_{12}^{\prime\prime\prime})-\lambda(G_{12}^{\prime\prime})$
$+\lambda(G_{12}^{\prime\prime})-\lambda(G^{\prime\prime})+\lambda(G^{\prime\prime})-\lambda(G)$
$\geq$ $\,\left[2b-\frac{k-5+6/k}{r-1}+2a^{\prime}+2-\frac{2}{r}+\frac{4(k-1)}{kr}+2(a-1)\right]\frac{1}{n}+O(n^{-2})$
$\geq$ $\,\left[-\frac{k-5+6/k}{r-1}+2-\frac{2}{r}+\frac{4(k-1)}{kr}\right]\frac{1}{n}+O(n^{-2})$
Now, the assumption $Q<c_{1}(r)$ and $1\leq k\leq 1/(1-Qr)$ imply that $02-\frac{k-5+6/k}{r-1}-\frac{4(k-1)}{kr}>$ 0. Since $k>4/3$, we have $\frac{4(k-1)}{kr}-\frac{2}{r}>-\frac{4(k-1)}{kr}$ and so the bracketed expression above is positive, giving a contradiction. ∎
This establishes that $|w_{i}-w_{j}|\leq 1$ for every $i,j$. We now complete the proof of Theorem 2.
Proof of Theorem 2.
As stated above, the assumption $Q<c_{1}(r)$ gives $02-\frac{k-5+6/k}{r-1}-\frac{4(k-1)}{kr}>$ 0, and we will use this estimate below. Let $H\in\mathrm{EX}(n,\mathcal{F})$, with partition $V_{1}\sqcup\cdots\sqcup V_{r}$ according to 1, and $v_{i}:=|V_{i}|$. If $v_{1}<v_{i}$ for some $i$ then $e(H_{i1})=e(H)$, so we may assume without loss of generality that $v_{1}\geq v_{i}$ for every $i$. Then either $w_{1}=v_{1}$ or $w_{1}=v_{1}-1$. Assume for a contradiction that $e(H)>e(G)$.
Case 1: $w_{1}=v_{1}$.
Subcase 1.1: $k\in\{1,2,3\}$. Then $H$ can be obtained from $G$ by a bounded number of edge deletions and additions, and by 41 we are done.
Subcase 1.2: $k\geq 4$. There is a graph $H^{\prime}$ be obtained from $H$ by a bounded number of edge modifications and a graph $G^{\prime}$ be obtained from $G$ by a bounded number of edge modifications such that $H^{\prime}$ and $G^{\prime}$ can be obtained from $T_{n,r}$ by embedding the same number of trees of order $k$ in the large part. Let $b=e(H)-e(H^{\prime})=O(1)$, $a=e(G)-e(G^{\prime})<b$. Then 43 gives $\lambda(H^{\prime})\geq\lambda(G^{\prime})-\left[\frac{k-5+6/k}{r-1}\right]\frac{1}{n}+O(n^{-2})$. Hence, using 41
$\lambda(H)-\lambda(G)$ $=\lambda(H)-\lambda(H^{\prime})+\lambda(H^{\prime})-\lambda(G^{\prime})+\lambda(G^{\prime})-\lambda(G)$
$\geq\left[2b-\frac{k-5+6/k}{r-1}-2a\right]\frac{1}{n}+O(n^{-2})$
$\geq\left[2-\frac{k-5+6/k}{r-1}\right]\frac{1}{n}+O(n^{-2}).$
Since $02-\frac{k-5+6/k}{r-1}>$ 0, this gives a contradiction.
Case 2: $w_{1}=v_{1}-1.$ Then without loss of generality, $w_{2}=w_{1}+1=v_{1}$. Let $G^{\prime}$ be obtained from $G$ by replacing every vertex in $W^{m}$ with an ordinary vertex. Let $a=e(G^{\prime})-e(G)=O(1)$.
Claim 18.
We have $\lambda(G_{21}^{\prime})-\lambda(G^{\prime})=-\left[\frac{4(k-1)}{kr}\right]\frac{1}{n}+O(n^{-2}).$
Proof.
Let $u^{\prime}$ be the ordinary vertex in $W_{2}$ which was transferred to obtain $G_{21}^{\prime}$. Let $\mu=\lambda(G^{\prime}),\rho=\lambda(G_{21}^{\prime})$, with corresponding eigenvectors $y,z$, scaled so that $y_{1}:=\frac{1}{\mu}\sum_{v\not\in W_{1}}y_{v}=1$ and $z_{1}:=\frac{1}{\rho}\sum_{v\not\in W_{1}\cup\{u^{\prime}\}}z_{v}=1$. Let $w_{2},\ldots,w_{h}=w_{1}+1$ and $w_{h+1},\ldots,w_{r}=w_{1}$. Similarly to 44, we obtain the following preliminary estimates. First, $\rho-\mu=O(n^{-1})$. For $v\in W_{1}$,
$y_{v}$ $=1+\frac{d_{W_{1}}(v)}{\mu}+\frac{d_{W_{1}}^{\prime}(v)}{\mu^{2}}+O(n^{-3})$
$z_{v}$ $=1+\frac{d_{W_{1}}(v)}{\mu}+\frac{d_{W_{1}}^{\prime}(v)}{\mu^{2}}+O(n^{-3}).$
Also $z_{u^{\prime}}=1$. Let $Y_{1}=\sum_{v\in W_{1}}y_{v},$ $Z_{1}=\sum_{W_{1}\cup\{u^{\prime}\}}z_{v}$. Then
$Y_{1}$ $=w_{1}+\frac{2e(W_{1})}{\mu}+\frac{\sum_{v\in W_{1}}d_{W_{1}}^{\prime}(v)}{\mu^{2}}+O(n^{2})$
$=w_{1}+\frac{2(k-1)}{k(r-1)}+O(n^{-1})$
$Z_{1}$ $=Y_{1}+1+O(n^{-2}).$
If $v\in W_{i}$ for $i\geq 2$, then
$y_{v}=\begin{cases}1+\frac{Y_{1}-w_{1}-1}{\mu+w_{1}}-\frac{Y_{1}-w_{1}-1}{(\mu+w_{1})^{2}}+O(n^{-3}),&2\leq i\leq h\\ 1+\frac{Y_{1}-w_{1}}{\mu+w_{1}},&h+1\leq i\leq r.\end{cases}$
For $v\in W_{2}-\{u^{\prime}\}$ or $v\in W_{i}$ with $i\geq 3$,
$z_{v}=\begin{cases}1+\frac{Y_{1}-w_{1}}{\mu+w_{1}}-\frac{Y_{1}-w_{1}}{(\mu+w_{1})^{2}}+O(n^{-3}),&3\leq i\leq h\\ 1+\frac{Y_{1}-w_{1}+1}{\mu+w_{1}}+O(n^{-3}),&i=2\text{ or }h+1\leq i\leq r.\end{cases}$
So,
$\mu=$ $\,(h-1)(w_{1}+1)\left(1+\frac{Y_{1}-w_{1}-1}{\mu+w_{1}}-\frac{Y_{1}-w_{1}-1}{(\mu+w_{1})^{2}}\right)$
$+(r-h)w_{1}\left(1+\frac{Y_{1}-w_{1}}{\mu+w_{1}}\right)+O(n^{-2})$
$\rho=$ $\,(h-2)(w_{1}+1)\left(1+\frac{Y_{1}-w_{1}}{\mu+w_{1}}-\frac{Y_{1}-w_{1}}{(\mu+w_{1})^{2}}\right)$
$+(r-h+1)w_{1}\left(1+\frac{Y_{1}-w_{1}+1}{\mu+w_{1}}\right)+O(n^{-2}).$
We obtain the initial estimate
$\mu=(r-1)w_{1}+h-1+\frac{2(k-1)}{kr}-\frac{h}{r}+\frac{1}{r}+O(n^{-1}).$
Then,
$\rho-\mu=$ $\,-1+\frac{(w_{1}+1)(w_{1}-Y_{1}+h-1)}{\mu+w_{1}}$
$-\frac{(w_{1}+1)(w_{1}-Y_{1}+h-1)}{(\mu+w_{1})^{2}}+\frac{w_{1}(Y_{1}-w_{1}+r-h+1)}{\mu+w_{1}}+O(n^{-2})$
$=$ $\,-1+\frac{rw_{1}}{\mu+w_{1}}+\frac{w_{1}-Y_{1}+h-1}{\mu+w_{1}}-\frac{w_{1}(w_{1}-Y_{1}+h-1)}{(\mu+w_{1})^{2}}+O(n^{-2}).$
Now
$\frac{rw_{1}}{\mu+w_{1}}$ $=1-\left[h-1+\frac{2(k-1)}{kr}-\frac{h}{r}+\frac{1}{r}\right]\frac{1}{n}+O(n^{-2})$
$\frac{w_{1}-Y_{1}+h-1}{\mu+w_{1}}$ $=\left[-\frac{2(k-1)}{k(r-1)}+h-1\right]\frac{1}{n}+O(n^{-2})$
$\frac{w_{1}(w_{1}-Y_{1}+h-1)}{(\mu+w_{1})^{2}}$ $=\left[-\frac{2(k-1)}{kr(r-1)}+\frac{h}{r}-\frac{1}{r}\right]\frac{1}{n}+O(n^{-2}).$
Thus,
$\rho-\mu=-\left[\frac{4(k-1)}{kr}\right]\frac{1}{n}+O(n^{-2}).$
Subcase 2.1: $k\in\{1,2,3\}$. Then $H$ can be obtained from $G_{21}^{\prime}$ by a bounded number of edge modifications. Let $b=e(H)-e(G_{21}^{\prime})=e(H)-e(G^{\prime})=e(H)-e(G)-a$, so $b+a\geq 1$. Then by 41 and 18, we have
$\lambda(H)-\lambda(G)$ $=\lambda(H)-\lambda(G_{21}^{\prime})+\lambda(G_{21}^{\prime})-\lambda(G^{\prime})+\lambda(G^{\prime})-\lambda(G)$
$\geq\left[2b-\frac{4(k-1)}{kr}+2a\right]\frac{1}{n}+O(n^{-2})>0,$
a contradiction.
Subcase 2.2: $k\geq 4$. Let $H^{\prime}$ be obtained from $H$ by deleting some $b=O(1)$ edges, and $G_{21}^{\prime\prime}$ be obtained from $G_{21}^{\prime}$ by deleting some $a^{\prime}=O(1)$ edges, so that $H^{\prime}$ and $G_{21}^{\prime\prime}$ are each obtained by embedding the same number of trees of order $k$ in a large part of $T_{n,r}$. Note that $e(H)>e(G)$ gives $b+a-a^{\prime}\geq 1$. Then by 41, 43, and 18, we have
$\lambda(H)-\lambda(G)=$ $\,\lambda(H)-\lambda(H^{\prime})+\lambda(H^{\prime})-\lambda(G_{21}^{\prime\prime})+\lambda(G_{21}^{\prime\prime})-\lambda(G_{21}^{\prime})$
$+\lambda(G_{21}^{\prime})-\lambda(G^{\prime})+\lambda(G^{\prime})-\lambda(G)$
$\geq\left[2b-\frac{k-5+6/k}{r-1}-2a^{\prime}-\frac{4(k-1)}{kr}+2a\right]\frac{1}{n}+O(n^{-2})$
$\geq\left[2-\frac{k-5+6/k}{r-1}-\frac{4(k-1)}{kr}\right]\frac{1}{n}+O(n^{-2}).$
The bracketed expression is positive, which gives a contradiction. ∎
Observe that in Subcase 2.2 of the proof of Theorem 2, if $e(H)=e(G)+1$, $k\geq 4$, and all the trees embedded in $H$ are paths while all the trees embedded in $G$ are stars, then according to 43, the equality
$\lambda(H)-\lambda(G)=\left[2-\frac{k-5+6/k}{r-1}-\frac{4(k-1)}{kr}\right]\frac{1}{n}+O(n^{-2})$
holds. In the next section we will show that there is a forbidden family $\mathcal{F}$ such that some $H\in\mathrm{EX}(n,\mathcal{F})$ and some $\mathcal{F}$-free graph $G$ in fact satisfy these conditions, and then use this to prove 1.
3 Counterexamples
3.1 Large extremal number
Proof of 1.
Suppose that $r\geq 3$, $n\equiv 1\pmod{r}$ is large enough, and $\lfloor n/r\rfloor=(n-1)/r\equiv 0\pmod{k}$ (note there are infinitely many such $n$). Let $G$ be obtained by embedding $U(K_{1,k-1},\lfloor n/r\rfloor)$ in a small part of $T_{n,r}$ and let $H$ be obtained by embedding $U(P_{k},\lceil n/r\rceil)$ in a large part of $T_{n,r}$ and then extending one of the paths by a single edge within the part, so that $e(H)=e(G)+1$. Let $V_{1}\sqcup\cdots\sqcup V_{r}$ and $U_{1}\sqcup\cdots\sqcup U_{r}$ be the corresponding partitions of $V(G)$ and $V(H)$, respectively, so that the trees are embedded in $V_{1}$ and $U_{1}$. Since
$k>\frac{r-1}{2}\left(2+\frac{5}{r-1}-\frac{4}{r}+\sqrt{\left(2+\frac{5}{r-1}-\frac{4}{r}\right)^{2}-\frac{4}{r-1}\left(\frac{6}{r-1}-\frac{4}{r}\right)}\right)$
we have
$\frac{1}{r-1}k+\frac{4}{r}-\frac{5}{r-1}+\left(\frac{6}{r-1}-\frac{4}{r}\right)k^{-1}>2.$
and in particular $k\geq 4$. Thus, by the observation below the proof of Theorem 2, this implies that $\lambda(H)<\lambda(G)$. So, to prove 1 it suffices to construct a family $\mathcal{F}$ such that $\mathrm{EX}(n,\mathcal{F})=\{H\}$ and $G$ is $\mathcal{F}$-free. Let $m>\max\{4,k-1\}$ and let $\mathcal{F}$ be the union of the following graph families.
  • (1)
    $T_{k}+T_{m(r-1),r-1}$ for every tree $T_{k}$ on $k$ vertices for which $T_{k}\not\simeq K_{1,k-1},P_{k}$
  • (2)
    $T_{k+1}+T_{m(r-1),r-1}$ for every tree $T_{k+1}$ on $k+1$ vertices for which $T_{k+1}\not\simeq P_{k+1}$
  • (3)
    $(2\cdot P_{k+1})+T_{m(r-1),r-1}$
  • (4)
    $(K_{1,k-1}\cup P_{k+1})+T_{m(r-1),r-1}$
  • (5)
    $C_{\ell}+T_{m(r-1),r-1}$ for $13\leq\ell\leq k+$ 1
  • (6)
    $K_{r+2}$
Claim 19.
$H$ is $\mathcal{F}$-free.
Proof.
Consider any copy $K$ of $T_{m(r-1),r-1}$ which appears in $H$ and assume for a contradiction that a forbidden subgraph of the form $F+K$ appears. If $V(K)\subseteq V-U_{1}$, then $N(V(K))=U_{1}$, so no forbidden graph of the form $F+K$ appears, by the definition of $H[U_{1}]$. Otherwise, $V(K)\ni v\in U_{1}$. If $|V(K)\cap U_{1}|\geq 3$, then $N(V(K)\cap U_{1})\cap U_{1}=\emptyset$, hence $V(F)\cap U_{1}=\emptyset$. Then $V(F)$ intersects two different parts $U_{2},U_{3}$, implying that $K$ is contained in the $r-2$ parts $U_{1},U_{4},\ldots,U_{r}$. But one can check $|V(K)\cap U_{i}|\leq m$ since $\Delta(H[U_{i}])\leq 2<m$ so $|V(K)|\leq(r-2)m$, a contradiction. Thus, $|V(K)\cap U_{1}|\leq 2$, so $|V(K)-U_{1}|\geq m(r-1)-2$, and thus $V(K)\cap U_{i}\neq\emptyset$ for each $i$ as $m(r-2)+2<m(r-1)-2$. Hence, $N(V(K))\subseteq U_{1}$, and the previous argument applies. This covers forbidden subgraphs of forms (1) through (5). Finally, if $H$ contains a copy $K$ of $K_{r+2}$, then $|V(K)\cap U_{i}|\leq 1$ for each $i>1$ and $|V(K)\cap U_{1}|\leq 2$. Hence $|V(K)|\leq r+1$, a contradiction. ∎
Claim 20.
$G$ is $\mathcal{F}$-free.
Proof.
Consider any copy $K$ of $T_{m(r-1),r-1}$ which appears in $G$. If $V(K)\subseteq V-V_{1}$, then $N(V(K))=V_{1}$, so no forbidden graph of the form $F+K$ appears, by the definition of $G[V_{1}]$. Otherwise, $V(K)\ni v\in V_{1}$. If $|V(K)\cap V_{1}|\geq k$, then $N(V(K)\cap V_{1})\cap V_{1}=\emptyset$, hence $V(F)\cap V_{1}=\emptyset$. Then $V(F)$ intersects two different parts $V_{2},V_{3}$, implying that $K$ is contained in the $r-2$ parts $V_{1},V_{4},\ldots,V_{r}$. But one can check $|V(K)\cap V_{i}|\leq m$ since $\Delta(G[V_{i}])\leq k-1<m$ so $|V(K)|\leq(r-2)m$, a contradiction. Thus, $|V(K)\cap V_{1}|\leq k-1$, so $|V(K)-V_{1}|\geq m(r-1)-k+1$, and thus $V(K)\cap V_{i}\neq\emptyset$ for each $i$ since $m\geq k-1$. Hence $N(V(K))\subseteq V_{1}$, and the previous argument applies. This covers forbidden subgraphs of forms (1) through (5). Finally, if $G$ contains a copy $K$ of $K_{r+2}$, then $|V(K)\cap U_{i}|\leq 1$ for each $i>1$ and $|V(K)\cap V_{1}|\leq 2$. Hence $|V(K)|\leq r+1$, a contradiction. ∎
Claim 21.
If $n$ is large enough we have $\mathrm{EX}(n,\mathcal{F})=\{H\}$.
Proof.
Let $H^{\prime}\in\mathrm{EX}(n,\mathcal{F})$. Following subsection 2.1 through 15 shows that $H^{\prime}$ is obtained from $K_{W_{1},\ldots,W_{r}}$ by embedding trees in $W_{1}$, where $|w_{i}-w_{j}|\leq 1$ for $i,j>1$ and $w_{1}-2\leq w_{i}\leq w_{1}+1$ for $i>1$. Suppose there are $h$ parts of size $w_{1}-2$ and $r-h-1$ parts of size $w_{1}-1$, so that
$w_{1}+h(w_{1}-2)+(r-h-1)(w_{1}-1)=n\Longrightarrow w_{1}=\frac{n-1}{r}+1+\frac{h}{r}.$
Now $n\equiv 1\pmod{r}$, the right-hand side is an integer, and $10\leq h\leq r-$ 1, so we have $h=0$, and in fact $K_{W_{1},\ldots,W_{r}}=T_{n,r}$. By (1) through (5), $H^{\prime}[W_{1}]$ is a disjoint union of trees, each of which is $P_{k+1}$, $P_{k}$, $K_{1,{k-1}}$, or of order smaller than $k$. Moreover, at most one tree is $P_{k+1}$ and such a tree can only exist if there is no $K_{1,k-1}$. From these restrictions and the fact that $\lfloor n/r\rfloor\equiv 0\pmod{k}$ it is clear that $e(H^{\prime}[W_{1}])$ (hence $e(H^{\prime})$) is maximized exactly when $H^{\prime}\simeq H$. ∎
By the discussion above, this completes the proof of 1. ∎
3.2 Chromatic number 3
Proof of 2.
Let $\mathcal{F}$ consist of $K_{4}$ and all graphs of the form $T_{4}+(m\cdot K_{1})$, where $T_{4}$ is a tree on four vertices. (Here $m$ is a sufficiently large constant.) Now $\chi(F)=3$ and $\mathcal{F}$ satisfies the assumptions of 1 with $Q=\frac{1}{3}$. Let $n=2p$, where $p\equiv 1\pmod{3}$. Let $G$ be the graph obtained by embedding $(p-1)/3\cdot P_{3}$ in the smaller part of $K_{p-1,p+1}$; let $H$ be the graph obtained by embedding $(p-1)/3\cdot P_{3}$ in one part of $K_{p,p}$, let $H^{\prime}$ be the graph obtained by embedding $((p-4)/3\cdot P_{3})\cup(2\cdot P_{2})$ in one part of $K_{p,p}$, and let $H^{\prime\prime}$ be the graph obtained by embedding $((p-1)/3\cdot P_{3})\cup P_{2}$ in the large part of $K_{p+1,p-1}$. Similarly to 21 one can prove $\mathrm{EX}(n,\mathcal{F})=\{H,H^{\prime},H^{\prime\prime}\}$. Now $G,H,H^{\prime},H^{\prime\prime}$ have equitable partitions with quotient matrices
$B=\begin{pmatrix}0&2&p+1\\ 1&0&p+1\\ \frac{p-1}{3}&\frac{2(p-1)}{3}&0\end{pmatrix},$ $\ \ C=\begin{pmatrix}0&2&0&p\\ 1&0&0&p\\ 0&0&0&p\\ \frac{p-1}{3}&\frac{2(p-1)}{3}&1&0\end{pmatrix},$
$C^{\prime}=\begin{pmatrix}0&2&0&p\\ 1&0&0&p\\ 0&0&1&p\\ \frac{p-4}{3}&\frac{2(p-4)}{3}&4&0\end{pmatrix},$ $\ \ C^{\prime\prime}=\begin{pmatrix}0&2&0&p\\ 1&0&0&p\\ 0&0&1&p\\ \frac{p-4}{3}&\frac{2(p-4)}{3}&4&0\end{pmatrix},$
respectively. We find that $\lambda(H),\lambda(H^{\prime}),\lambda(H^{\prime\prime})<p+\frac{2}{3}-\frac{1}{5p}<\lambda(G)$ by substituting this value into the characteristic polynomials. Since $G\not\in\mathrm{EX}(n,\mathcal{F})$, this implies that $\mathrm{SPEX}(n,\mathcal{F})\cap\mathrm{EX}(n,\mathcal{F})=\emptyset$. ∎
4 Proof of 1
Since
$\frac{r-1}{2}\left(2+\frac{5}{r-1}-\frac{4}{r}+\sqrt{\left(2+\frac{5}{r-1}-\frac{4}{r}\right)^{2}-\frac{4}{r-1}\left(\frac{6}{r-1}-\frac{4}{r}\right)}\right)$
is not an integer, it is strictly less than its ceiling $k$. Hence, 1 gives that $c(r)\leq\frac{k-1}{kr}$. Suppose $Q<\frac{k-1}{kr}$. Let $\mathcal{F}$ be any finite family of graphs with $\chi(\mathcal{F})=r+1$ and $\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,r})+Qn$ for large enough $n$. Then we have
$k^{\prime}$ $:=k(\mathcal{F})\leq\left\lfloor\frac{1}{1-Qr}\right\rfloor\leq k-1$
$<\frac{r-1}{2}\left(2+\frac{5}{r-1}-\frac{4}{r}+\sqrt{\left(2+\frac{5}{r-1}-\frac{4}{r}\right)^{2}-\frac{4}{r-1}\left(\frac{6}{r-1}-\frac{4}{r}\right)}\right)$
and so $\frac{k^{\prime}-1}{k^{\prime}r}<c_{1}(r)$. Moreover, 1 gives that $\mathrm{ex}(n,\mathcal{F})\leq e(T_{n,r})+\frac{k^{\prime}-1}{k^{\prime}r}n+O(1)$ for $n$ large enough, as $k^{\prime}=k(\mathcal{F})$. Hence, we may apply Theorem 2 to show that $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$ for $n$ large enough. Therefore, $c(r)\geq Q$.∎
5 Concluding remarks
Theorem 2 gives a condition on the value of $\mathrm{ex}(n,\mathcal{F})$ which guarantees that $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$, when $\chi(\mathcal{F})=r+1\geq 4$. This condition is sharp for many values of $r$. As shown by 2, Theorem 2 does not hold if we allow $r=2$ and $k(\mathcal{F})=3$. We attempted to provide a similar counterexample in the case of $r=2,k(\mathcal{F})=2$ but were unable to do so because $\lambda(H)$ is in fact greater than $\lambda(G)$ by a term of $\Theta(n^{-2})$, where $G$ and $H$ are graphs defined analogously to the ones in subsection 3.2. This fact, combined with the result of [wang2023conjecture], suggested 1. A solution to 1 would be interesting as it potentially requires stronger techniques than those of this paper and those of [wang2023conjecture].
We believe that the methods of this paper are general enough to apply to many specific families $\mathcal{F}$ which do not satisfy the assumptions of Theorem 2. Even for $\mathcal{F}$ such that $\mathrm{SPEX}(n,\mathcal{F})\not\subseteq\mathrm{EX}(n,\mathcal{F})$, large parts of the proof of Theorem 2 go through and may complete much of the work in determining the extremal graphs.
Besides 1, the open questions we are most interested in are the following.
  • (1)
    Can the assumption that $\mathcal{F}$ is finite be dropped from Theorem 2?
  • (2)
    What is the smallest value of $|\mathcal{F}|$ for which an example proving 1 can be constructed?
[…]